### v1

# Contrastive Learning on Multimodal Analysis of Electronic Health Records

Abstract
Electronic health record (EHR) systems contain a wealth of multimodal clinical data including structured data like clinical codes and unstructured data such as clinical notes. However, many existing EHR-focused studies has traditionally either concentrated on an individual modality or merged different modalities in a rather rudimentary fashion. This approach often results in the perception of structured and unstructured data as separate entities, neglecting the inherent synergy between them. Specifically, the two important modalities contain clinically relevant, inextricably linked and complementary health information. A more complete picture of a patient’s medical history is captured by the joint analysis of the two modalities of data. Despite the great success of multimodal contrastive learning on vision-language, its potential remains under-explored in the realm of multimodal EHR, particularly in terms of its theoretical understanding. To accommodate the statistical analysis of multimodal EHR data, in this paper, we propose a novel multimodal feature embedding generative model and design a multimodal contrastive loss to obtain the multimodal EHR feature representation. Our theoretical analysis demonstrates the effectiveness of multimodal learning compared to single-modality learning and connects the solution of the loss function to the singular value decomposition of a pointwise mutual information matrix. This connection paves the way for a privacy-preserving algorithm tailored for multimodal EHR feature representation learning. Simulation studies show that the proposed algorithm performs well under a variety of configurations. We further validate the clinical utility of the proposed algorithm in real-world EHR data.
Keywords: Natural language processing, textual data, structured data, representation learning, singular value decomposition.
1 Introduction
The growing accessibility of Electronic Health Record (EHR) data presents numerous opportunities for clinical research, ranging from patient profiling (Halpern et al.,, 2016) to predicting medical events (Choi et al.,, 2017). However, the complexity increases with the multimodal nature of EHR data, which encompasses diverse clinical data from patient demographics and genetic information to unstructured textual data like clinical notes, and structured data such as diagnostic and procedure codes, medication orders, and lab results.
A key challenge in EHR-focused research lies in effectively merging these different data types and ensuring that their clinical aspects are meaningfully represented. Research has shown the benefits of integrating structured and unstructured data for tasks like automated clinical code assignment (Scheurwegs et al.,, 2016), managing chronic diseases (Sheikhalishahi et al.,, 2019), and pharmacovigilance (Stang et al.,, 2010). While these different modalities serve as complementary data sources, there is significant overlap and correlation among these data (Qiao et al.,, 2019). Joint representation of both structured and narrative data into a more manageable low-dimensional space where similar features are grouped closely can significantly improve the utility of both data types. This representation learning technique has gained popularity for its ability to capture and represent the intricate relationships among various EHR features.
Despite the extensive research on EHR feature representation, most existing studies have primarily focused on either structured (Choi et al., 2016a, ; Kartchner et al.,, 2017; Hong et al.,, 2021; Zhou et al.,, 2022) or unstructured data modalities (De Vine et al.,, 2014; Choi et al., 2016b, ; Beam et al.,, 2019; Alsentzer et al.,, 2019; Huang et al.,, 2020; Lehman and Johnson,, 2023). For instance, Alsentzer et al., (2019) adapted the BERT model (Devlin et al.,, 2019) to the clinical domain by training on the MIMIC-III clinical notes (Johnson et al.,, 2016). It was extremely time/resource-consuming to train such a black-box model. De Vine et al., (2014) aligned free-text features with the Unified Medical Language System (UMLS)’s common concept unique identifier (CUI) space (McInnes et al.,, 2007). They then utilized the skip-gram algorithm (Mikolov et al.,, 2013) on concatenated concept documents to learn the CUI embeddings. Similarly, Choi et al., 2016b and Beam et al., (2019) employed singular value decomposition (SVD) on a pointwise mutual information (PMI) matrix derived from CUI co-occurrences in unstructured text to generate CUI embeddings. This approach was also adopted by Hong et al., (2021) for code embeddings. This SVD-PMI algorithm not only ensures scalability but also upholds data privacy through the use of aggregate co-occurrence data, offering a scalable variant of the skip-gram algorithm (Levy and Goldberg,, 2014). Its interpretability is further highlighted in the dynamic log-linear topic model developed by Arora et al., (2016).
Recent studies have emerged on leveraging multimodal EHR features for enhanced predictive modeling, as demonstrated by the work of Khadanga et al., (2019); Zhang et al., (2020); Bardak and Tan, (2021); Gan et al., (2023). Specifically, Gan et al., (2023) enhanced code and CUI embeddings by employing the SVD-PMI algorithm, facilitating the integration of structured and unstructured data. Despite these advancements, methods relying on straightforward data merging may not fully account for the intricate interactions within multimodal data, potentially leading to biases. This issue will be further examined in our theoretical (Section 3) and numerical analyses (Sections 4 and 5).
Addressing these limitations, Liu et al., (2022) introduced a multimodal pre-trained language model incorporating a cross-attention mechanism to enrich EHR representations across structured and unstructured data domains. Additionally, the domain has seen progress in multimodal contrastive learning strategies, as evidenced by research from Li and Gao, (2022); Yin et al., (2023); Wang et al., (2023). These strategies, drawing inspiration from successful vision-language models like the Contrastive Language-Image Pre-Training (CLIP) by Radford et al., (2021), aim to train unified representations of diverse data modalities. However, these approaches, grounded in deep neural networks, face challenges due to their “black-box” nature, including a lack of theoretical underpinning, computational complexity, and concerns over privacy since they necessitate access to individual patient data. These factors contribute to their limited applicability in the context of EHR data.
While some theoretical analyses of multimodal learning exist, their applicability to EHR data has been limited. Groundbreaking studies like Huang et al., (2021) have illustrated the benefits of multimodal learning, showing that learning across multiple modalities can reduce population risk compared to single-modality methods. Recently, Deng et al., (2023) theoretically proved the zero-shot transfer ability of CLIP. Furthermore, Nakada et al., (2023) explored multimodal contrastive learning’s performance under a spiked covariance model. However, these studies do not directly apply to the unique discrete feature structure of EHR data, leaving an unaddressed theoretical gap in understanding multimodal contrastive learning’s application in healthcare. Bridging this gap is vital, as it lays a solid foundation for multimodal contrastive learning’s development and implementation in healthcare, maximizing its potential to improve patient care and drive medical research forward.
To overcome these limitations, we introduce the Contrastive Learning Algorithm for Integrated Multimodal Electronic health records (CLAIME). Our findings confirm that CLAIME is not only an effective tool for deriving multimodal EHR feature representations but also a method that respects privacy by requiring only aggregated data. Additionally, we propose a novel multimodal feature embedding generative model (2.1) in Section 2.1, designed to enhance statistical analysis of multimodal EHR data. This model is notable for its interpretability and accurate portrayal of EHR data generation. It distinguishes itself from earlier word vector generative models (Arora et al.,, 2016, 2018; Lu et al.,, 2023; Xu et al.,, 2022) by (1) enabling the integration of multimodal EHR features, and (2) allowing patient heterogeneity by incorporating error terms specific to patients, thus increasing the model’s robustness. Within this generative framework, we validate the consistency of the CLAIME algorithm and clarify the relationship between multimodal feature embeddings and a multimodal PMI matrix. The proposed algorithm is also privacy-preserving since it only requires summary-level data, opening doors for collaboration across multiple institutions. Our research also fills a theoretical void in the analysis of multimodal contrastive learning for EHR data.
The rest of the paper is structured as follows. Section 2 elaborates on the proposed method. Section 3 presents the theoretical properties of our algorithm. In Section 4, simulation studies demonstrate the algorithm’s effectiveness under various configurations. Section 5 further validates our algorithm’s clinical applicability in EHR studies. Finally, Section 6 includes discussions.
2 Method
2.1 Notation
For any matrix $\mathbf{A}$, let $\|\mathbf{A}\|$, $\|\mathbf{A}\|_{\max}$ and $\|\mathbf{A}\|_{\mathrm{F}}$ be its operator norm, entrywise maximum norm and Frobenius norm, respectively. We define $\mathbf{P}_{p}(\mathbf{A})$ as the top-$p$ right singular vectors of $\mathbf{A}$. When the right singular vectors are not unique, we choose arbitrary singular vectors. Denote $s_{j}(\mathbf{A})$ as the $j$-th largest singular value of $A$. Let $\mathcal{O}_{d,p}$ ($d\geq p$) denote a set of $d\times p$ orthonormal matrices. For two sequences of positive numbers $\{a_{k}\}$ and $\{b_{k}\}$, we write $a_{k}\lesssim b_{k}$ or $a_{k}=O(b_{k})$ or $b_{k}\gtrsim a_{k}$ or $b_{k}=\Omega(a_{k})$ if there exists some constant $C>0$ such that $a_{k}\leq Cb_{k}$ for all $k$. We denote $a_{k}\ll b_{k}$ or $a_{k}=o(b_{k})$ if $\lim_{k\rightarrow\infty}a_{k}/b_{k}=0$. For any positive integer $I$, let $[I]=\{1,2,\cdots,I\}$. We write $a\vee b$ and $a\wedge b$ to denote $\max(a,b)$ and $\min(a,b)$, respectively. We denote by $\bm{e}_{j}$ the $j$-th unit vector in $\mathbb{R}^{d}$ where its $j$-th element is one, and all other elements are zero.
2.2 Model Assumptions
Assume that the collections of codes from the structured modality and CUIs from the unstructured modality are denoted by ${\cal W}^{(1)}:=[d_{1}]$ and ${\cal W}^{(2)}:=[d]\setminus[d_{1}]$ respectively, where $d=d_{1}+d_{2}$, with $d$ representing the total number of unique features across both modalities. Suppose that we have $n$ independent patients. For each patient $i$, the observed codes and CUIs are denoted as $\{w^{(1)}_{i,t}\}_{t\in[T_{i}^{(1)}]}$ and $\{w^{(2)}_{i,t}\}_{t\in[T_{i}^{(2)}]}$, where $w^{(1)}_{i,t}\in{\cal W}^{(1)}$ and $w^{(2)}_{i,t}\in{\cal W}^{(2)}$. The sizes of these sets are given by $T_{i}^{(1)}\geq 2$ and $T_{i}^{(2)}\geq 2$, respectively. We model the probability of observing specific codes and CUIs for the $i$-th patient based on their embeddings $\mathbf{v}_{w}^{\star}\in\mathbb{R}^{p}$ as follows:
$\mathbb{P}(w^{(1)}_{i,t}=w|\bm{c}_{i},{\bm{\epsilon}}_{i}^{(1)})$ $=\frac{{\exp}(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w})}{\sum_{w^{\prime}\in{\cal W}^{(1)}}{\exp}(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}})},\ \ w\in{\cal W}^{(1)},t\in[T_{i}^{(1)}],\text{ and }$ (2.1)
$\mathbb{P}(w^{(2)}_{i,t}=w|\bm{c}_{i},{\bm{\epsilon}}_{i}^{(2)})$ $=\frac{{\exp}(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w})}{\sum_{w^{\prime}\in{\cal W}^{(2)}}{\exp}(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}})},\ \ w\in{\cal W}^{(2)},t\in[T_{i}^{(2)}]\,.$
Here $\bm{c}_{i}\sim N(\bm{0},\mathbf{I}_{p})$ represents a latent vector specific to patient $i$, reflecting their clinical state. The error terms, ${\bm{\epsilon}}^{(1)}_{i}=(\epsilon_{i,1},\dots,\epsilon_{i,d_{1}})^{\top}\sim N(\bm{0},{\bm{\Sigma}}_{1})$ and ${\bm{\epsilon}}^{(2)}_{i}=(\epsilon_{i,d_{1}+1},\dots,\epsilon_{i,d})^{\top}\sim N(\bm{0},{\bm{\Sigma}}_{2})$, account for patient-specific variations and address the variability not captured by the term $\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle$, where ${\bm{\Sigma}}_{1}$ and ${\bm{\Sigma}}_{2}$ are some unknown positive semi-definite matrices.
2.3 CLAIME Algorithm
We define code and CUI embedding matrices as $\mathbf{V}_{1}^{\star}=(\mathbf{v}_{1}^{\star},\ldots,\mathbf{v}_{d_{1}}^{\star})^{\top}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}^{\star}=(\mathbf{v}_{d_{1}+1}^{\star},\ldots,\mathbf{v}_{d}^{\star})^{\top}\in\mathbb{R}^{d_{2}\times p}$, respectively, and aim to infer $\mathbf{V}^{\star}=\big{(}\mathbf{V}_{1}^{\star\top},\mathbf{V}_{2}^{\star\top}\big{)}^{\top}\in\mathbb{R}^{d\times p}$. The embeddings should reflect clinical semantics, meaning that highly similar (e.g. rheumatoid arthritis and juvenile rheumatoid arthritis) or related (e.g. fasting glucose and type II diabetes) EHR entities should have close embeddings. Before introducing our algorithm, we first define aggregate co-occurrence matrices $\mathbf{C}^{(M,M)}$ and $\mathbf{D}^{(M,M^{\prime})}$ for $M,M^{\prime}\in\{1,2\}$ across different modalities as:
$\mathbf{C}^{(M,M)}(w,w^{\prime})=\sum_{i=1}^{n}\mathbf{C}^{(M,M)}_{i,i}(w,w^{\prime}),\ \ \mathbf{D}^{(M,M^{\prime})}(w,w^{\prime})=\sum_{i=1}^{n}\mathbf{D}^{(M,M^{\prime})}_{i,i}(w,w^{\prime})$
where $\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M)}]:t\neq s,w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ and $\mathbf{D}^{(M,M^{\prime})}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M^{\prime})}]:w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ for $i,j\in[n]$ and $M\in\{1,2\}$. Further, we define the marginal co-occurrence of $w\in{\cal W}^{(M)}$ as:
$\gamma_{w}^{(M)}=\mathbf{C}^{(M,M)}(w,\cdot)=\sum_{w^{\prime}\in{\cal W}^{(M)}}\mathbf{C}^{(M,M)}(w,w^{\prime}).$ (2.2)
We introduce $S_{q}^{(M)}=(n^{-1}\sum_{i=1}^{n}(T_{i}^{(M)})^{q})^{1/q}$ for $q\geq 1$ and $S_{1}^{(1,2)}=n^{-1}\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}$ as scaling factors. CLAIME utilizes the multimodal contrastive learning loss defined as:
$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})=\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}\sum_{i\neq j}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{j,s}}\rangle}{\gamma_{w_{i,t}}^{(1)}\gamma_{w_{j,s}}^{(2)}}$ (2.3)
$\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{1}{nS_{1}^{(1,2)}}\sum_{i=1}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{i}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{i,s}}\rangle}{\gamma_{w_{i,t}}^{(1)}\gamma_{w_{i,s}}^{(2)}}+\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\top}\|_{\mathrm{F}}^{2}.$
Here $\mathbf{V}_{1}=(\mathbf{v}_{1},\ldots,\mathbf{v}_{d_{1}})^{\top}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}=(\mathbf{v}_{d_{1}+1},\ldots,\mathbf{v}_{d})^{\top}\in\mathbb{R}^{d_{2}\times p}$, $\lambda>0$ serves as a regularization coefficient, $\gamma_{w}^{(M)},M=1,2$ are weights chosen based on the frequency of $w$, as defined in (2.2). Our theoretical analysis in Section 3 motivates the choice of $\gamma_{w}^{(M)}$ to guide the minimizer towards $\mathbf{V}^{\star}$. The essence of the multimodal contrastive learning loss in CLAIME is to enhance the representation of similar features across different modalities by bringing them closer together while distancing those that are dissimilar. In the context of EHR data, this translates to aligning the embeddings of codes and CUIs that have clinical correlations and separating those that do not. The aim is to maximize the inner product of embeddings for features that co-occur within the same patient’s data.
Remark 2.1.
Our CLAIME framework can be naturally extended to non-lienar loss functions. Note that our CLAIME loss function (2.3) can be written as
$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=-\frac{1}{\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}}\sum_{i=1}^{n}\quantity{T_{i}^{(1)}T_{i}^{(2)}s_{ii}-\alpha_{T}\sum_{j:j\neq i}^{n}T_{i}^{(1)}T_{j}^{(2)}s_{ij}}+R(\mathbf{V}_{1},\mathbf{V}_{2}),$
where $R(\mathbf{V}_{1},\mathbf{V}_{2})$ is a smooth regularizer, and
$\alpha_{T}=\frac{\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}}{\sum_{i\neq j}^{n}T_{i}^{(1)}T_{j}^{(2)}},\ \ s_{ij}=\frac{1}{T_{i}^{(1)}T_{j}^{(2)}}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{j,s}}\rangle}{\gamma_{w_{i,t}}^{(1)}\gamma_{w_{j,s}}^{(2)}},\ \ \ \ \ \ i,j\in[n].$
One can consider the following non-linear loss function analogous to the CLIP loss function (Radford et al.,, 2021):
$\mathcal{L}_{\textnormal{CLAIME}}^{\prime}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=-\frac{1}{\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}}\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}\log\frac{\exp(s_{ii}/\tau)}{\sum_{j:j\neq i}^{n}T_{i}^{(1)}T_{j}^{(2)}\exp(s_{ij}/\tau)}+R(\mathbf{V}_{1},\mathbf{V}_{2}),$
which becomes equivalent to the loss (2.3) when $\tau\to\infty$.
To obtain $\mathbf{\widehat{V}}=(\mathbf{\widehat{V}}_{1}^{\top},\mathbf{\widehat{V}}_{2}^{\top})^{\top}=\arg\min\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ efficiently in practice, we note that $\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ can be expressed in terms of pair-wise co-occurrences of concepts as:
$\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}^{\star}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{j\neq i}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{j,s}=w^{\prime})$ (2.4)
$\quad-\frac{1}{nS_{1}^{(1,2)}}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}^{\star}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{i}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{i,s}=w^{\prime})+\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\top}\|_{\mathrm{F}}^{2}\,.$
Subsequently, via arguments given in Supplementary S2, we have the following proposition.
Proposition 2.1.
We have
$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=\frac{\lambda}{2}\Big{\|}\mathbf{V}_{1}\mathbf{V}_{2}^{\top}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant}),$
where $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}=\{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})\}_{w\in{\cal W}^{(1)},w^{\prime}\in{\cal W}^{(2)}}$ with
$\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ $:=\frac{\mathbf{C}^{(1,1)}(\cdot,\cdot)\mathbf{C}^{(2,2)}(\cdot,\cdot)}{\mathbf{C}^{(1,1)}(w,\cdot)\mathbf{C}^{(2,2)}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{D}^{(1,2)}(w,w^{\prime})}{\mathbf{D}^{(1,2)}(\cdot,\cdot)}-\frac{\mathbf{C}^{(c)}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}),$
and $\mathbf{C}^{(c)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j\neq i}^{n}\mathbf{D}^{(1,2)}_{i,j}(w,w^{\prime}).$
Proposition 2.1 shows that $\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ can be related to the SVD of an empirical association matrix, where its element $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ estimates the association between the features $w$ and $w^{\prime}$. We will later demonstrate the convergence of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ to the population PMI matrix in Section 3. As a result, $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ can be viewed as a modified estimator of the population PMI matrix. The final embeddings $\mathbf{\widehat{V}}_{1}\mathbf{\widehat{V}}_{2}^{\top}$ are inferred through a rank-$p$ SVD of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, preserving data privacy and offering a scalable estimation. To be more specific, let $\widehat{\mathbf{U}}_{1}\widehat{\bm{\Lambda}}\widehat{\mathbf{U}}_{2}$ denote the rank-$p$ SVD of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, where $\widehat{\mathbf{U}}_{1}\in\mathbb{R}^{d_{1}\times p}$ and $\widehat{\mathbf{U}}_{2}\in\mathbb{R}^{d_{2}\times p}$ are the matrices of left and right singular vectors, respectively, and $\widehat{\bm{\Lambda}}\in\mathbb{R}^{p\times p}$ is a diagonal matrix with its diagonal elements being the top $p$ singular values. Then, we set $\mathbf{\widehat{V}}_{1}=\widehat{\mathbf{U}}_{1}\widehat{\bm{\Lambda}}^{1/2}$ and $\mathbf{\widehat{V}}_{2}=\widehat{\mathbf{U}}_{2}\widehat{\bm{\Lambda}}^{1/2}$. It is worth noting that the selection of the regularization parameter $\lambda$ does not play a crucial role, and for the sake of simplicity, we will assign it a value of $1$ in our upcoming numerical analyses.
2.4 Comparison between CLAIME and Simple Concatenation
We next contrast CLAIME with the simple approach of ignoring between-modality differences between different modalities. Dealing with multimodal data often presents difficulties, leading to conventional methods that overlook the differences between various modalities. A basic strategy commonly adopted is to simply merge the two modalities through direct concatenation, after which algorithms initially intended for unimodal data are applied. However, this rudimentary treatment of multimodal data may lead to substantial bias due to the inherent heterogeneity between different modalities. To illustrate, consider the concatenated data for the $i$-th patient represented as
$\{w_{i,t}\}_{t\in[T_{i}]}=\Big{(}w_{i,1}^{(1)},\dots,w_{i,T_{i}^{(1)}}^{(1)},w_{i,1}^{(2)},\dots,w_{i,T_{i}^{(2)}}^{(2)}\Big{)},\text{ where }T_{i}=T_{i}^{(1)}+T_{i}^{(2)}\,.$
A popular method to handle such data is the SVD-PMI algorithm, as referenced in (Levy and Goldberg,, 2014; Gan et al.,, 2023). In this context, we establish the co-occurrence matrices for the concatenated dataset as follows:
$\mathbf{C}=\begin{bmatrix}\mathbf{C}^{(1,1)}&\mathbf{D}^{(1,2)}\\ \mathbf{D}^{(2,1)}&\mathbf{C}^{(2,2)}\end{bmatrix},$
where $\mathbf{C}^{(M,M)}$ and $\mathbf{D}^{(M,M^{\prime})}$ are defined in Section 2.3. Subsequently, the empirical concatenated PMI matrix, denoted as $\widehat{\mathbb{PMI}}=\{\widehat{\mathbb{PMI}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$, is formulated as
$\widehat{\mathbb{PMI}}(w,w^{\prime})$ $=\log\frac{\mathbf{C}(w,w^{\prime})\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)},$ (2.5)
where $\mathbf{C}(w,\cdot)=\sum_{w^{\prime}=1}^{d}\mathbf{C}(w,w^{\prime})$ and $\mathbf{C}(\cdot,\cdot)=\sum_{w=1}^{d}\mathbf{C}(w,\cdot)$. Following this, we conduct a rank-$p$ eigen-decomposition of $\widehat{\mathbb{PMI}}$, represented as $\mathbf{\widehat{U}}_{\textnormal{Con}}\widehat{\bm{\Lambda}}_{\textnormal{Con}}\mathbf{\widehat{U}}_{\textnormal{Con}}$. The estimator of $\mathbf{V}^{\star}$ is then achieved by setting $\mathbf{\widehat{V}}_{\textnormal{Con}}=\mathbf{\widehat{U}}_{\textnormal{Con}}\widehat{\bm{\Lambda}}_{\textnormal{Con}}^{1/2}$. We refer to this method as “Concate”.
The second prevalent technique is contrastive learning (CL), applied directly to the concatenated dataset. Specifically, we define the contrastive loss for $\mathbf{V}\in\mathbb{R}^{d\times p}$ as follows:
$\mathcal{L}_{\textnormal{CL}}(\mathbf{V})=-\frac{1}{\sum_{i=1}^{n}T_{i}(T_{i}-1)}\sum_{i=1}^{n}\sum_{t\in[T_{i}]}\sum_{s\in[T_{i}]\setminus\{t\}}\frac{\langle\mathbf{v}_{w_{i,t}},\mathbf{v}_{w_{i,s}}\rangle}{\gamma_{w_{i,t}}\gamma_{w_{i,s}}}$ (2.6)
$\quad\quad+\frac{1}{\sum_{i=1}^{n}\sum_{j\neq i}^{n}T_{i}T_{j}}\sum_{i\neq j}\sum_{t\in[T_{i}]}\sum_{s\in[T_{j}]}\frac{\langle\mathbf{v}_{w_{i,t}},\mathbf{v}_{w_{j,s}}\rangle}{\gamma_{w_{i,t}}\gamma_{w_{j,s}}}+\frac{\lambda}{2}\|\mathbf{V}\mathbf{V}^{\top}\|_{\mathrm{F}}^{2},$
where $\gamma_{w}=\mathbf{C}(w,\cdot)$. Let $\widehat{\mathbf{V}}_{\textnormal{CL}}=\arg\min\mathcal{L}_{\textnormal{CL}}(\mathbf{V})$. Similar to Proposition 2.1, we have the following proposition. Here, we define
$\mathbf{D}=\begin{bmatrix}\mathbf{C}^{(1)}&\mathbf{C}^{(c)}\\ \mathbf{C}^{(c)\top}&\mathbf{C}^{(2)}\end{bmatrix}\text{ with }\mathbf{C}^{(M)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j\neq i}^{n}\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})\text{ for }w,w^{\prime}\in{\cal W}^{(M)}.$
Proposition 2.2.
We have
$\mathcal{L}_{\textnormal{CL}}(\mathbf{V})$ $=\frac{\lambda}{2}\Big{\|}\mathbf{V}\mathbf{V}^{\top}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CL}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant}),$
where $\widehat{\mathbb{PMI}}_{\textnormal{CL}}=\{\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$ with
$\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime}):=\frac{\mathbf{C}(\cdot,\cdot)\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{C}(w,w^{\prime})}{\mathbf{C}(\cdot,\cdot)}-\frac{\mathbf{D}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}).$
Proposition 2.2 reveals that obtaining $\mathbf{\widehat{V}}_{\textnormal{CL}}$ by minimizing the loss $\mathcal{L}_{\textnormal{CL}}$ is essentially equivalent to applying PCA on the matrix $\widehat{\mathbb{PMI}}_{\textnormal{CL}}$. The proof for Proposition 2.2 can be found in Supplementary S4.4. This method is referred to as CL. Nevertheless, as we will demonstrate in Section 3, both $\mathbf{\widehat{V}}_{\textnormal{Con}}$ and $\mathbf{\widehat{V}}_{\textnormal{CL}}$ are not optimal solutions.
3 Theoretical Analysis
In this section, we investigate the theoretical properties of CLAIME and draw comparisons with the Concate and CL methods. First, let us establish some fundamental concepts. For $w\in{\cal W}^{(M)}$, and $t\in[T_{i}^{(M)}]$, denote $X_{i,w}^{(M)}(t)=\mathbb{I}\{w^{(M)}_{i,t}=w\}$. According to the model (2.1), conditioned on $\bm{c}_{i}$ and ${\bm{\epsilon}}_{i}^{(M)}$, the variable $X_{i,w}^{(M)}(t)$ follows a Bernoulli distribution, with the probability of occurrence given by
$p_{i,w}^{(M)}=\mathbb{E}[X_{i,w}^{(M)}(t)|\bm{c}_{i},{\bm{\epsilon}}_{i}^{(M)}]=\frac{\exp(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w})}{\sum_{w^{\prime}\in{\cal W}^{(M)}}\exp(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}})}\,.$ (3.1)
Here, the probability $p_{i,w}^{(M)}$ is a function of the discourse vector $\bm{c}_{i}$ and the noise term ${\bm{\epsilon}}_{i}^{(M)}$, and is independent of $t$. Next, we denote the expected value of $p_{i,w}^{(M)}$ as $p_{w}^{(M)}=\mathbb{E}[p_{i,w}^{(M)}]$. Further, we define the co-occurrence of a feature $w\in{\cal W}^{(M)}$ at time $t$ for the $i$-th patient and feature $w^{\prime}\in{\cal W}^{(M^{\prime})}$ at time $s$ for the $j$-th patient as $X_{i,j,w,w^{\prime}}^{(M,M^{\prime})}(t,s)=\mathbb{I}\{w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}$ for $t\neq s$. This variable follows a Bernoulli distribution, with the probability of occurrence denoted as $p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}$. Furthermore, we define the expected value for the case of $i=j$ as $p_{w,w^{\prime}}^{(M,M^{\prime})}=\mathbb{E}[p_{i,i,w,w^{\prime}}^{(M,M^{\prime})}]$. It is important to note that when $i\neq j$, the expected value $\mathbb{E}[p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}]$ equals the product $p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}$. Now, we introduce the population PMI matrix $\mathbb{PMI}$ defined as
$\mathbb{PMI}:=\begin{pmatrix}\mathbb{PMI}^{(1,1)}&\mathbb{PMI}^{(1,2)}\\ \mathbb{PMI}^{(2,1)}&\mathbb{PMI}^{(2,2)}\end{pmatrix}\text{ with }\mathbb{PMI}^{(M,M^{\prime})}(w,w^{\prime})=\log\frac{p_{w,w^{\prime}}^{(M,M^{\prime})}}{p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}}$ (3.4)
for $w\in{\cal W}^{(M)}$, $w^{\prime}\in{\cal W}^{(M^{\prime})}$ with $M,M^{\prime}\in\{1,2\}$. Before proceeding with our theoretical analysis, let us establish some general assumptions. Based on our data generation model, we can only identify the left singular space and the singular values of $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$. Thus, without loss of generality, we can assume that the singular value decomposition of $\mathbf{V}_{M}^{\star}$ is $\mathbf{V}_{M}^{\star}=\mathbf{U}_{M}^{\star}\bm{\Lambda}_{M}^{\star}$ for $M\in\{1,2\}$, where $\mathbf{U}_{M}^{\star}\in\mathcal{O}_{d_{M},p}$ represents the left singular vectors, and $\bm{\Lambda}_{M}^{\star}\in\mathbb{R}^{p\times p}$ is a diagonal matrix containing the corresponding singular values. The embedding can only be identified up to a mean shift. For simplicity, we assume $\mathbf{V}_{M}^{\star\top}\bm{1}_{d_{M}}=\bm{0}$.
Assumption 3.1.
For $M\in\{1,2\}$, assume that $\|\mathbf{U}_{M}^{\star}\|_{2,\infty}\lesssim\sqrt{p/d_{M}}$, $\|\bm{\Lambda}_{M}^{\star}\|/s_{p}(\bm{\Lambda}_{M}^{\star})\lesssim 1$, and $\|\bm{\Lambda}_{M}^{\star}\|\ll 1$.
Assumption 3.1 is widely known as the incoherence constant condition in the literature appearing in matrix completion (Candès and Recht,, 2009), PCA (Zhang et al.,, 2022), and the analysis of representation learning algorithm (Ji et al.,, 2021).
We introduce the notation $\underline{d}=\min(d_{1},d_{2})$, and ${\bm{\Sigma}}=\mathrm{diag}({\bm{\Sigma}}_{1},{\bm{\Sigma}}_{2})$.
Assumption 3.2.
Assume that $\log d\ll d_{1}\wedge d_{2}$, $p\log^{12}d\ll d_{1}\wedge d_{2}$, $p\ll(\log\log d)^{2}$, $n\gg p^{2}d^{5}\log^{2}d$ and $p^{3}\log^{6}d\ll\underline{d}^{3/2}/d^{1/2}$.
Assumption 3.2 imposes a condition on the dimension of the low-rank representation. This assumption is technically necessary in bounding the normalizing constant of the log-linear word production model (Lemma S5.2). Although, we can relax the assumption of $p\ll(\log\log d)^{2}$ into $p\ll\log^{2}d$, we stick to the current assumption for the sake of clarity.
Assumption 3.3.
Assume that $\|\mathrm{diag}({\bm{\Sigma}}_{M})\|_{\max}\lesssim p/d_{M}$ for $M\in\{1,2\}$ and $1/p\ll s_{p}({\bm{\Sigma}})\leq\|{\bm{\Sigma}}_{1}\|\vee\|{\bm{\Sigma}}_{2}\|\lesssim 1$.
Assumption 3.3 implies that the signal-to-noise ratio is bounded above and below. Note that the signal, measured by the smallest positive singular value of $(1/p)\mathbf{V}_{M}^{\star}\mathbf{V}_{M^{\prime}}^{\star\top}$ for $M,M^{\prime}\in\{1,2\}$ is of order $1/p$ from Assumption 3.1.
Theorem 3.1.
Under Assumptions 3.1, 3.2 and 3.3, we have
$\Big{\|}\mathbb{PMI}^{(M,M)}-\quantity(\frac{1}{p}\mathbf{V}_{M}^{\star}\mathbf{V}_{M}^{\star\top}+{\bm{\Sigma}}_{M})\Big{\|}$ $\lesssim\frac{p^{3/2}}{d_{M}}\log^{6}d,\ \ M\in\{1,2\}$
$\Big{\|}\mathbb{PMI}^{(1,2)}-\frac{1}{p}\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star\top}\Big{\|}$ $\lesssim\frac{p^{3/2}}{\underline{d}}\log^{6}d,$
and hence
$\Big{\|}\mathbb{PMI}-\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}})\Big{\|}$ $\lesssim\frac{p^{2}}{\underline{d}}\log^{6}d.$
The proof of Theorem 3.1 is presented in Supplementary S4.1. This theorem demonstrates that the PMI matrix for each modality can be closely approximated by the embedding space plus the noise covariance matrix ${\bm{\Sigma}}_{M}$. When the noise level ${\bm{\Sigma}}_{M}$ is substantial, the PMI value $\mathbb{PMI}^{(M,M)}(w,w^{\prime})$ for a single modality may significantly deviate from $\langle\mathbf{v}_{w}^{\star},\mathbf{v}_{w^{\prime}}^{\star}\rangle/p$. On the other hand, the cross-modal PMI matrix $\mathbb{PMI}^{(1,2)}$ can be directly approximated by $\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star\top}/p$. The theorem below affirms that $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an effective estimator for $\mathbb{PMI}^{(1,2)}$.
Theorem 3.2.
Suppose that $n\gg d^{2}\log^{2}d$. Under Assumptions 3.1-3.3 and 3.5, we have
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}-\mathbb{PMI}^{(1,2)}}_{\max}$ $\lesssim\frac{d_{1}d_{2}}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}^{3/2}}\log^{6}d\ \ \text{ and }$
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}-\mathbb{PMI}^{(1,2)}}$ $\lesssim\frac{\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}}\log^{6}d$
with probability $1-\exp(-\Omega(\log^{2}d))$.
The proof of Theorem 3.2 is given in Supplementary S4.3. By Assumption 3.2, it implies that $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an asymptotically unbiased estimate of $\mathbb{PMI}^{(1,2)}$. Combining Theorems 3.1 and 3.2, we get the rate of CLAIME. We also provide a rate of learning dual representations jointly by CLAIME. Namely, we consider concatenating the left singular vectors of two matrices $\mathbf{\widehat{V}}_{1}$ and $\mathbf{\widehat{V}}_{2}$. Since $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\top})$ are only identifiable up to rotation, the subspace spanned by the column vectors of $\mathbf{P}_{p}([\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{2})]$ depends on the choice of rotations. Hence we introduce a matrix $\mathbf{H}$, and define the estimated joint representation by $\mathbf{\widehat{U}}_{\mathbf{H}}:=\mathbf{P}_{p}([\mathbf{H}\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{2}^{\top})])$, where $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\top})$ and $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\top})$ are arbitrary top-$p$ left singular vectors of $\mathbf{\widehat{V}}_{1}$ and $\mathbf{\widehat{V}}_{2}$, respectively. We then measure the best performance of $\mathbf{\widehat{U}}_{\mathbf{H}}$ with respect to $\mathbf{H}$ by $\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$, where $\mathbf{U}^{\star}\in\mathcal{O}_{d,p}$ is the concatenated embedding vectors defined as $\mathbf{U}^{\star}:=\mathbf{P}_{p}([\mathbf{U}_{1}^{\star\top},\mathbf{U}_{2}^{\star\top}])$.
Theorem 3.3.
Suppose that $n\gg d^{2}\log^{2}d$. Under Assumptions 3.1-3.3 and 3.5,
$\norm{\mathbf{\widehat{V}}_{1}\mathbf{\widehat{V}}_{2}^{\top}-\frac{1}{\lambda p}\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star\top}}\lesssim\frac{1}{\lambda}\quantity(\frac{\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}}\log^{6}d),$
$\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{U}_{1}^{\star})\|_{\mathrm{F}}\vee\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\top}),\mathbf{U}_{2}^{\star})\|_{\mathrm{F}}\lesssim\frac{p^{3/2}\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d$
and
$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\frac{p^{3/2}\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d$
with probability $1-\exp(-\Omega(\log^{2}d))$.
The proof of Theorem 3.3 is presented in Supplementary S4.3. On the contrary, we can show that the empirical concatenated PMI matrix $\widehat{\mathbb{PMI}}$ and $\widehat{\mathbb{PMI}}_{\textnormal{CL}}$ derived from the CL loss (2.6) are biased estimators for $\mathbb{PMI}$.
Theorem 3.4.
Suppose that $n\gg d^{2}\log^{2}d$. Under Assumptions 3.1-3.3 and 3.5, we have
$\norm{\widehat{\mathbb{PMI}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{Con}})}_{\max}\lesssim\frac{d^{2}}{\sqrt{n}}\log d,\quad\norm{\widehat{\mathbb{PMI}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{Con}})}\lesssim\frac{d^{5/2}}{\sqrt{n}}\log d,$
and
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CL}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{CL}})}_{\max}$ $\lesssim\frac{d^{2}}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}^{3/2}}\log^{6}d,$
$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CL}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{CL}})}$ $\lesssim\frac{d^{5/2}}{\sqrt{n}}\log d+\quantity(\frac{d}{\underline{d}})^{1/2}\frac{p^{3/2}}{\underline{d}}\log^{6}d.$
Here the matrices $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ reflect the bias terms defined in Supplementary S4.2 and S4.4.
Remark 3.1.
The bias matrices $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ depend only on $T_{i}^{(M)}$’s, which can be effectively removed since $T_{i}^{(M)}$’s are directly observed.
The proof of Theorem 3.4 is presented in Supplementary S4.4. Based on this theorem, we are ready to show the rate of $\widehat{\mathbf{V}}_{\textnormal{CL}}$ and $\mathbf{\widehat{V}}_{\textnormal{Con}}$. Before that, we need the following assumption.
Assumption 3.4.
Assume that
$\frac{s_{p}({\bm{\Sigma}})-s_{p+1}({\bm{\Sigma}})}{s_{p}({\bm{\Sigma}})}\gtrsim 1.$
Assumption 3.4 is known as the eigengap condition for the top-$p$ singular vectors of ${\bm{\Sigma}}$. A similar assumption has been used in the analysis of contrastive learning (Ji et al.,, 2021). This enables us to identify the top-$p$ singular vectors of $(1/p)\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}}$, resulting in the following theorem.
Theorem 3.5.
Suppose that $n\gg d^{2}\log^{2}d$ and $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1/p$. Under Assumptions 3.1-3.5,
$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{\textnormal{Con}}^{\top}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{Con}}))}_{\mathrm{F}}$ $\lesssim\frac{p^{3/2}d^{5/2}}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d,$
$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\widehat{\mathbf{V}}_{\textnormal{CL}}^{\top}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{CL}}))}_{\mathrm{F}}$ $\lesssim\frac{p^{3}d^{5/2}}{\sqrt{n}}\log d+\quantity(\frac{d}{\underline{d}})^{1/2}\frac{p^{9/2}}{\underline{d}}\log^{6}d$
holds with probability $1-\exp(-\Omega(\log^{2}d))$.
The proof Theorem 3.5 is given in Supplementary S4.4. It shows that both Concate and CL will yield biased estimators for the embeddings $\mathbf{V}^{\star}$ due to the inherent bias terms and covariance matrix of the noise. While the bias terms $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ can be removed, the influence of ${\bm{\Sigma}}$ remains irreducible due to its unobserved nature. Comparing Theorems 3.3 and 3.5, we find that CLAIME achieves a sharper rate than Concate and CL even when $\mathbf{B}_{\textnormal{CL}}$, $\mathbf{B}_{\textnormal{Con}}$ and ${\bm{\Sigma}}$ go to zero. As a result, CLAIME is a better estimator than Concate and CL.
To quantify the effect of $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$, we define the following empirical moments of $T_{i}^{(M)}$: $S_{q}^{(M)}:=((1/n)\sum_{i\in[n]}T_{i}^{(M)q})^{1/q}$. We also define the empirical cross-moments as $S_{q}^{(1,2)}:=((1/n)\sum_{i\in[n]}(T_{i}^{(1)})^{q/2}(T_{i}^{(2)})^{q/2})^{1/q}$ for $q=2,4$. Note that when the length of patient data is the same, i.e., $T_{i}^{(M)}=T^{(M)}$ for all $i\in[n]$ for $M=1,2$, $S_{q}^{(M)}=T^{(M)}$ for all $q\geq 1$.
Assumption 3.5.
For $M=1,2$, assume that $1\gtrsim S_{1}^{(1)}S_{1}^{(2)}/S_{2}^{(1,2)2}\gg 1/n$, $S_{1}^{(M)2}/S_{2}^{(M)2}\gg 1/n$ and $S_{2}^{(M)2}\gg S_{1}^{(M)}$. Also assume that $S_{1}^{(M)}\gtrsim S_{2}^{(M)}\gtrsim S_{4}^{(M)}$ and $S_{2}^{(1,2)}\gtrsim S_{4}^{(1,2)}$.
Assumption 3.5 requires that the empirical variance of $T_{i}^{(M)}$ dominates the mean of $T_{i}^{(M)}$, and their moments are comparable to each other. For the same length setting $T_{i}^{(M)}\equiv T^{(M)}$ for all $i\in[n]$, the assumption is satisfied if $T^{(M)}\to\infty$ and $n\to\infty$. For the bias term $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$, we have the following lemma. The proof of Lemma 3.1 is given in Supplementary S4.7.
Lemma 3.1.
Suppose that Assumption 3.5 holds. If $T_{i}^{(1)}=T_{i}^{(2)}$, then,
$\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim d\frac{S_{1}^{(1)}}{S_{2}^{(1)2}}.$
Furthermore, if $T_{i}^{(1)}=T_{i}^{(2)}\equiv T$ for all $i\in[n]$ for some $T>0$, then
$\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim\frac{d}{T}.$
We then characterize the effect of ${\bm{\Sigma}}$. We will show that when ${\bm{\Sigma}}$ satisfies suitable conditions, the two existing methods Concate and CL will not be an ideal estimator of $\mathbf{V}^{\star}$.
Assumption 3.6.
Assume $\|\sin\Theta(\mathbf{U}_{1}^{\star},\mathbf{P}_{p}({\bm{\Sigma}}_{1}))\|_{\mathrm{F}}\wedge\|\sin\Theta(\mathbf{U}_{2}^{\star},\mathbf{P}_{p}({\bm{\Sigma}}_{2}))\|_{\mathrm{F}}\geq\sqrt{(2/3)p}$.
Assumption 3.6 holds if the column space of $\mathbf{P}_{p}({\bm{\Sigma}}_{1})$ is nearly orthogonal to the column space of $\mathbf{U}^{\star}_{1}$ and that of $\mathbf{U}^{\star}_{2}$ nearly orthogonal to that of $\mathbf{P}_{p}({\bm{\Sigma}}_{2})$. Roughly speaking, this assumption requires that we can well separate the features and noises based on their directions in each modality.
Theorem 3.6.
Suppose that Assumptions 3.1-3.6 hold. If $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1$, then,
$\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{Con}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\gtrsim\sqrt{p}\ \ \text{ and }\ \ \|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{CL}},\mathbf{U}^{\star})\|_{\mathrm{F}}\gtrsim\sqrt{p}$
with probability $1-\exp(-\Omega(\log^{2}d))$.
The formal statement is available in Theorem S4.8. Theorem 3.6 shows that there is a huge deviation from the estimators of Concate and CL to the true parameters as $\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{Con}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ and $\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{CL}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ are both lower bounded by the order of $\sqrt{p}$. Since for any $\mathbf{\widehat{U}},\mathbf{U}^{\star}\in\mathcal{O}_{d,p}$, $\|\sin\Theta(\mathbf{\widehat{U}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ can be trivially upper bounded by $\sqrt{p}$, this result shows that $\mathbf{\widehat{U}}_{\textnormal{Con}}$ and $\mathbf{\widehat{U}}_{\textnormal{CL}}$ can hardly capture any signal in $\mathbf{U}^{\star}$.
We present an example construction of the error covariance matrix where we expect to see the performance difference among CLAIME, Concate, and CL.
Corollary 3.1.
Choose $\mathbf{P}^{(1)}\in\mathcal{O}_{d_{1},p}$ satisfying $[\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\in\mathcal{O}_{d_{1},2p}$ and $\|[\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\|_{2,\infty}^{2}\lesssim p/d_{1}$. Similarly choose $\mathbf{P}^{(2)}\in\mathcal{O}_{d_{2},p}$ such that $[\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\in\mathcal{O}_{d_{2},2p}$ and $\|[\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\|_{2,\infty}^{2}\lesssim p/d_{2}$. Let ${\bm{\Sigma}}_{1}:=\mathbf{P}^{(1)}\mathbf{P}^{(1)\top}$, ${\bm{\Sigma}}_{2}:=(1/2)\mathbf{P}^{(2)}\mathbf{P}^{(2)\top}$, $\mathbf{V}^{\star}_{1}=\mathbf{U}^{\star}_{1}$ and $\mathbf{V}^{\star}_{2}=\mathbf{U}^{\star}_{2}$. For these noise covariance matrices, Assumptions 3.1, 3.3 and 3.4 hold. In addition, Assumption 3.6 holds since $\mathbf{P}^{(1)\top}\mathbf{V}_{1}^{\star}=\mathbf{P}^{(2)\top}\mathbf{V}_{2}^{\star}=\mathbf{O}_{p\times p}$. Under additional assumptions on $(T_{i}^{(M)})_{i\in[n]}$, $d_{1}$ and $d_{2}$ (Assumptions 3.2 and 3.5), we obtain
$\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{Con}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\gtrsim\sqrt{p},\ \ \|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{CL}},\mathbf{U}^{\star})\|_{\mathrm{F}}\gtrsim\sqrt{p}\ \ \text{ and }$
$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\frac{p^{3/2}\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d.$
Corollary 3.1 directly follows from Theorems 3.3 and 3.6. Note that the factor $1/2$ in the definition of ${\bm{\Sigma}}_{2}$ guarantees the eigengap condition at the $p$-th largest singular value of the joint noise covariance matrix ${\bm{\Sigma}}$. The noise covariance in Corollary 3.1 implies that the subspace spanned by the noise is orthogonal to the subspace spanned by the signal, so that the noisy fluctuations in feature frequency distribution occurs simultaneously to many words, in a separable way from the signal. Under this condition, Concate and CL learn the eigenspace corresponding to noise covariance matrix, since those methods are vulnerable to large noise strength. On the other hand, CLAIME can effectively escape from the noise, and learns core representations when $n$ and $\bar{d}$ are sufficiently large.
4 Simulation Studies
In this section, we assess CLAIME’s performance and compare it to CL and Concate via simulation studies. When patient-level data is not available, these three methods can be deployed by performing SVD on different aggregated PMI matrices introduced in Section 2. Since patient-level data is also available in the synthesized datasets, we can additionally apply the gradient descent-based multimodal contrastive learning method (CLAIME-GD) with target function at (2.3), and the gradient descent-based contrastive learning method (CL-GD) with target function at (2.6), respectively. The suffix “-GD” is added to distinguish them from their SVD-based counterparts. To enhance training efficiency, only $10$ negative samples per patient are stochastically sampled. The penalty parameter $\lambda$ is set to $1$, while the learning rate starts at $10^{-4}$ and decays by a factor of $10$ every $10$ epochs to ensure convergence. Both gradient-based algorithms converge when their loss functions change by less than $10^{-6}$.
We generate patient embedding vector $\mathbf{c}_{i}\in\mathbb{R}^{p}$ for each patient $i$ from a multivariate normal distribution $N(\mathbf{0},\mathbf{I}_{p})$ independently. We construct the code and CUI embedding matrices $\mathbf{V}_{1}^{\star}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}^{\star}\in\mathbb{R}^{d_{2}\times p}$, respectively, by generating each entry from a standard normal distribution. The row means of $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$ are standardized to zero, and their maximum singular values are standardized to $1$. Then, we follow the data generation process in (2.1) to generate $T_{i}^{(1)}$ number of codes and $T_{i}^{(2)}$ number of CUIs for the $i$-th patient, where $T_{i}^{(1)},T_{i}^{(2)}$ follow a Poisson distribution with the mean of $50$ and only non-zero values are retained. Two different cases of error covariance structures are considered.
For Case 1, we set $\mathbf{\Sigma}^{(M)}=\text{diag}(\mathbf{I}_{p/2},\mathbf{0})/c\in\mathbb{R}^{d_{M}\times d_{M}}$ for $M=1$ or $2$, where $c>0$ is a constant controlling the signal-to-noise ratio (SNR) of the data generating process. For Case 2, we generated $\sigma_{w}\overset{i.i.d.}{\sim}\text{Unif}(0,1)$ and set ${\bm{\Sigma}}_{w,w^{\prime}}^{(M)}=\rho^{|w-w^{\prime}|}\sigma_{w}\sigma_{w^{\prime}}/2$ for $w,w^{\prime}\in[d_{M}]$ and $M\in\{1,2\}$, where $\rho\in(0,1)$ controls both the structure and the SNR of the data generating process. Specifically, a larger $\rho$ leads to a greater signal-to-noise ratio in the synthesized data.
For both cases, three experiments are separately conducted to compare the performance of CLAIME, CLAIME-GD, CL, CL-GD, and Concate under different sample sizes $n$, feature sizes $d$, and SNR. Their performance is evaluated using the metric $\text{Err}(\widehat{\mathbf{V}},\mathbf{V}^{\star})=\max_{M=1,2}\|\widehat{\mathbf{V}}_{M}\widehat{\mathbf{V}}_{M}^{\top}-\mathbf{V}_{M}^{\star}\mathbf{V}_{M}^{\star\top}\|_{\mathrm{F}}$, which measures the distance between the estimated and true embedding subspaces. We first fix the SNR by setting $c=0.2$ for Case 1 and $\rho=0.8$ for Case 2. In the first experiment, we vary the sample sizes $n$ in $\{2\times 10^{4},4\times 10^{4},6\times 10^{4},8\times 10^{4},10^{5}\}$, while fixing $d=2d_{1}=2d_{2}=1000$ and $p=4$. In the second experiment, we vary $d$ in $\{200,400,600,800,1000\}$ and set $d_{1}=d_{2}=d/2$, while fixing $n=10^{5}$ and $p=4$. For the third experiment, we first fix $n=10^{5}$, $d=2d_{1}=2d_{2}=1000$ and $p=4$. Then, we vary $c$ equally spaced between $0.2$ and $1$ for Case 1, and we vary $\rho$ equally spaced between 00 and $0.9$ for Case 2.
Figure 1: In the top panel, the error metric is plotted against varying $n$, $d$ and $c$ for Case 1. In the bottom panel, the error metric is plotted against varying $n$, $d$ and $\rho$ for Case 2.
The error metrics $\text{Err}(\widehat{\mathbf{V}},\mathbf{V}^{\star})$, averaged over $500$ repetitions, are presented in Figure 1. The results for CL and CL-GD, CLAIME and CLAIME-GD are very close to each other, as indicated by the proximity of their lines in Figures 1 (a)–(c). This suggests that employing gradient-based methods on patient-level data (i.e., CL-GD and CLAIME-GD) is comparable to utilizing SVD on summary-level PMI matrices (i.e., CL and CLAIME). This observation aligns with the theoretical findings in Propositions 2.1 and 2.2. On one hand, in Figures 1(a) and (b), the errors for CLAIME and CLAIME-GD decrease with an increase in $n$ or a decrease in $d$. This behavior is anticipated and can be theoretically inferred from the upper bounds provided in Theorems 3.13.3. In contrast, the errors for Concate, CL, and CL-GD remain unchanged. This can be theoretically justified by their lower bounds being irrelevant to both $n$ and $d$, as shown in Theorem 3.6. From the top panel of Figure 1(c), it can be observed that all methods exhibit larger errors when $c$ decreases, or in other words, the SNR decreases. Comparing to the other methods, CLAIME and CLAIME-GD demonstrate superior performance in the small SNR regions where $c\leq 0.5$. Similarly, the bottom panel of Figure 1(c) shows a consistent increase in errors across all methods with higher values of $\rho$. CLAIME and CLAIME-GD again significantly outperform the other methods when $\rho$ is larger than $0.5$. This supports our initial motivation that the proposed method will be more robust in cases when the data is noisier.
5 Application to Electronic Health Records Studies
We consider training a joint representation of codified and narrative CUI features using summary EHR data of patients with at least 1 diagnostic code of rhematoid arthritis (RA) at Mass General Brigham (MGB). The MGB RA EHR cohort consist of 53,716 patients whose longitudinal EHR data have been summarized as co-occurrence counts of EHR concepts within a $30$-day window. We include all EHR diagnosis, medication, and procedure codes which have been rolled up to higher concept levels: diagnostic codes to PheCodes [^1] https://phewascatalog.org/phecodes, procedure codes to clinical classification system (CCS) [^2] https://www.hcup-us.ahrq.gov/toolssoftware/ccs_svcsproc/ccssvcproc.jsp, medication codes to ingredient level RxNorm codes (Liu et al.,, 2005). We include all EHR codified features along with CUIs that have occurred more than $15$ times in the cohort. This results in $d=4,668$ features, including $d_{1}=3,477$ codified features with $1,776$ PheCodes for diseases, $238$ CCS codes for procedures, and $1,463$ RxNorm codes for medications, and $d_{2}=1,048$ CUI features.
Based on the summary-level co-occurrence matrix, we can derive estimators for CLAIME, CL, and Concate, respectively. To assess the quality of the obtained embeddings, we utilize a benchmark previously introduced in Gan et al., (2023). Specifically, we conduct an evaluation focusing on the embeddings’ capability to identify established relationships among EHR concept pairs. Various categories of known relations, including similarity and relatedness, have been curated from online knowledge sources. Similar pairs of codified concepts were mainly generated based on code hierarchies such as the PheCode hierarchy. Similar pairs of CUIs were extracted from relationships within the online clinical database UMLS. Additionally, we assessed the similarity between mapped code-CUI pairs, leveraging UMLS to map codified concepts to CUIs. For relatedness among CUI-CUI pairs, we extracted information from the UMLS, considering major classes like “may treat or may prevent”, “classifies”, “differential diagnosis”, “method of”, and “causative”. Furthermore, related code-CUI pairs and code-code pairs were curated by mapping disease CUIs to PheCodes, drugs to RxNorm, and procedures to CCS categories. These mapped code pairs were then utilized to evaluate the system’s ability to detect relatedness among codified features or the relatedness between codified and NLP features.
To assess different relationship types, we compute cosine similarities between embedding vectors of related pairs and randomly selected pairs. This approach enables the calculation of the Area Under the Curve (AUC) of the cosine similarities, providing a metric for distinguishing known pairs from random pairs. The random pairs selected have the same semantic type as the known pairs. For example, in investigating the relationship “may treat or may prevent,” we only consider disease-drug pairs. This evaluation strategy allows us to quantify and compare the performance of the EHR embeddings across various relationship types.
Pairs Type Group CLAIME CL Concate Num
CUI-Code Similar CUI-PheCode 0.920 0.920 0.920 345
CUI-RXNORM 0.972 0.919 0.918 49
summary 0.915 0.910 0.909 394
Related May Treat (Prevent) 0.775 0.753 0.753 1302
Classifies 0.905 0.895 0.895 744
ddx 0.773 0.765 0.766 1169
Causative 0.775 0.741 0.742 524
summary 0.800 0.783 0.784 3739
CUI-CUI Similar Parent 0.848 0.855 0.856 422
Sibling 0.796 0.763 0.766 865
summary 0.813 0.794 0.796 1287
Related May Treat (Prevent) 0.767 0.765 0.766 257
Classifies 0.936 0.924 0.925 45
ddx 0.741 0.771 0.773 95
Method_of 0.809 0.796 0.800 15
summary 0.756 0.752 0.752 421
Code-Code Similar PheCode Hierachy 0.914 0.909 0.909 3868
Related ddx 0.790 0.796 0.795 5819
Classifies 0.896 0.900 0.900 4525
May Treat (Prevent) 0.734 0.727 0.725 4598
Causative 0.754 0.756 0.754 2468
summary 0.798 0.799 0.798 17410
Table 1: AUCs of between-vector cosine similarity in detecting known similar or related pairs of codes, CUIs, and codes vs CUIs for embeddings trained by CLAIME, CL, and Concate methods. The last column indicates the number of known pairs.
The results are presented in Table 1. It can be observed that CLAIME excels in embedding relationships across two distinct modalities: code and CUIs. In particular, for the similar CUI-RxNorm group, CLAIME achieves about $6\%$ gain in AUC compared to CL and Concate. It is also much better at capturing causative relationships between codes and CUIs, achieving a $4.4\%$ gain in AUC. Moreover, CLAIME is slightly better in representing the parent-child hierarchical relationships between NLP concepts, with a 2% gain in AUC compared to the other two methods.
Since the data comes from the RA cohort, it is of interest for us to further perform case studies for RA-related codified and CUI concepts. We select $7$ key concepts for detailed analysis: rheumatoid arthritis, C-reactive proteins, stiffness of joint, Reiter’s disease, swollen joint, pneumococcal vaccine, and leflunomide.
For each of the $7$ concepts, we calculate its cosine similarity with all remaining concepts using the embeddings generated by CLAIME, Concate, and CL, respectively. Concepts are ranked by similarity, and a subset comprising the top-$100$ rankings from any one of the three methods is chosen as positive control. Then, we randomly select an equal number of other concepts as negative controls. This process produces an evaluation set for each of the $7$ concepts. Concordance between similarity scores (from CLAIME, Concate, CL) and relevance scores (from GPT3.5 (OpenAI,, 2023), GPT4 (Achiam et al.,, 2023) ) is assessed using Kendall’s tau rank correlation. Relevance scores are obtained by prompting GPT3.5 and GPT4 to rate concept relevance on a scale of 00 to $1$ between the given concept and its respective evaluation set.
Figure 2: Feature-specific Kendall’s ranking correlation between pairs of CLAIME, Concate and CL with GPT3.5 and GPT4 feature scores for 7 concepts, namely, rheumatoid arthritis (RA), C-reactive proteins (CP), stiffness of joint (SoJ), Reiter’s disease (RD), swollen joint (SJ), pneumococcal vaccine (PV), and leflunomide (LF). Note that “GPTx” refers to GPT3.5 on the left and GPT4 on the right, and the lines for “CL-GPTx” overlap with “Concate-GPTx” and hence cannot be seen from the figure.
The results, depicted in Figure 2, reveal that Concate and CL exhibit remarkably similar behavior, with the “CL-GPTx” lines obscured beneath “Concate-GPTx”. CLAIME can effectively encode and preserve relations between these medical concepts and their associated features. Notably, negative rank correlations can be observed for CL-GPT3.5/4 and Concate-GPT3.5/4 regarding “leflunomide” and “pneumococcal vaccine” related codes and CUIs. To gain more insights, some detailed examples are provided in Table 2.
ID Desc Concate CL CLAIME GPT3.5 GPT4
Relatedness or similarity with Leflunomide
PheCode:714.1 Rheumatoid arthritis 0.35 0.35 0.80 0.9 0.9
PheCode:714.2 Juvenile rheumatoid arthritis 0.36 0.37 0.78 0.9 0.8
C0242708 DMARDs 0.36 0.36 0.81 0.9 1.0
RXNORM:5487 Hydrochlorothiazide 0.80 0.81 0.01 0.1 0.2
RXNORM:6809 Metformin 0.79 0.79 0.01 0.2 0.2
RXNORM:2556 Citalopram 0.78 0.78 0.05 0.1 0.1
Relatedness or similarity with Pneumococcal vaccine
C0151735 Injection site reaction NOS 0.29 0.29 0.74 0.8 0.8
CCS:228 Prophylactic vaccinations and inoculations 0.29 0.29 0.64 0.9 1.0
C1445860 Protein antibody 0.04 0.04 0.61 0.6 0.7
RXNORM:10689 Tramadol 0.80 0.80 0.12 0.1 0.1
RXNORM:2556 Citalopram 0.75 0.75 0.15 0.1 0.1
RXNORM:11248 Cyanocobalamin 0.81 0.81 0.18 0.2 0.1
Table 2: Cosine similarity between leflunomide, pneumococcal vaccine and some selected features.
More specifically, leflunomide is a type of disease-modifying antirheumatic drugs (DMARDs) used to treat rheumatoid arthritis by reducing inflammation and permanent damage. Table 2 shows that both Concate and CL fail to capture the close connection between leflunomide, RA, and DMARDs. Additionally, they overestimate similarity with unrelated drugs like hydrochlorothiazide, metformin, and citalopram. In the case of pneumococcal vaccine, it is a vaccine targeting streptococcus pneumoniae infections and can potentially cause injection site reactions. It is related to protein antibodies as they all target the immune system (von Elten et al.,, 2014). Both Concate and CL struggle to capture this similarity and instead overestimate similarity with unrelated drugs like tramadol, cyanocobalamin, and citalopram, used for pain relief, vitamin B12 deficiency, and depressive disorders respectively.
6 Discussion and Conclusion
In this paper, we propose the noisy multimodal log-linear production model for analyzing the multimodal EHR data and present the privacy-preserving algorithm CLAIME to estimate the multimodal feature embeddings with theoretical justification. Our algorithm is readily available for federated learning when multiple healthcare systems want to co-train their model but can not share patient-level data. It is also of great interest to consider a setting where different systems have overlapping but non-identical feature sets, in which case a block-wise matrix completion algorithm may be useful (Zhou et al.,, 2021).
Although our current analysis focuses on two modalities, it can be extended to more than two modalities. The rich EHR data contains many different modalities such as genetic data and image data. It will be interesting to extend our current framework to accommodate these data. Besides, it will also be interesting to incorporate the patient demographic in the generative model to consider the effects of variables like sex and age.
This paper focuses on a linear loss function, while non-linear loss functions such as mentioned in Remark 2.1 also deserve to be explored. Another interesting problem is the estimation of the patient embedding $\bm{c}_{i}$. The patient embedding is useful for many downstream tasks such as identifying “patients like me”.

### v2

@@
 Abstract
-Electronic health record (EHR) systems contain a wealth of multimodal clinical data including structured data like clinical codes and unstructured data such as clinical notes.
-However, many existing EHR-focused studies has traditionally either concentrated on an individual modality or merged different modalities in a rather rudimentary fashion.
-This approach often results in the perception of structured and unstructured data as separate entities, neglecting the inherent synergy between them.
-Specifically, the two important modalities contain clinically relevant, inextricably linked and complementary health information.
-A more complete picture of a patient’s medical history is captured by the joint analysis of the two modalities of data.
+Electronic health record (EHR) systems capture a wealth of multimodal clinical data, encompassing both structured clinical codes and unstructured clinical notes.
+Yet, many EHR-focused studies have traditionally examined these modalities in isolation or combined them using simplistic methods.
+Such approaches reinforce the view of structured and unstructured data as separate, overlooking the intrinsic synergy between them.
+In reality, these modalities are deeply interconnected, each containing clinically relevant and complementary information that, when integrated effectively, can provide a more comprehensive understanding of patient health.
+Integrating these modalities enables a richer, more complete representation of a patient’s medical history, offering deeper insights into their health.
 Despite the great success of multimodal contrastive learning on vision-language, its potential remains under-explored in the realm of multimodal EHR, particularly in terms of its theoretical understanding.
-To accommodate the statistical analysis of multimodal EHR data, in this paper, we propose a novel multimodal feature embedding generative model and design a multimodal contrastive loss to obtain the multimodal EHR feature representation.
+To support statistical analysis of multimodal EHR data, we propose a novel multimodal feature embedding generative model and design a multimodal contrastive loss to obtain the multimodal EHR feature representation.
 Our theoretical analysis demonstrates the effectiveness of multimodal learning compared to single-modality learning and connects the solution of the loss function to the singular value decomposition of a pointwise mutual information matrix.
@@
 The growing accessibility of Electronic Health Record (EHR) data presents numerous opportunities for clinical research, ranging from patient profiling (Halpern et al., , 2016) to predicting medical events (Choi et al., , 2017).
-However, the complexity increases with the multimodal nature of EHR data, which encompasses diverse clinical data from patient demographics and genetic information to unstructured textual data like clinical notes, and structured data such as diagnostic and procedure codes, medication orders, and lab results.
-A key challenge in EHR-focused research lies in effectively merging these different data types and ensuring that their clinical aspects are meaningfully represented.
+However, the complexity increases with the multimodal nature of EHR data, which encompasses diverse data from patient demographics and genetic information to unstructured textual data like clinical notes, as well as structured data such as diagnostic and procedure codes, medication orders, and lab results.
+A key challenge in EHR-focused research lies in effectively merging these different data types and ensuring that their clinical aspects are meaningfully and coherently represented.
 Research has shown the benefits of integrating structured and unstructured data for tasks like automated clinical code assignment (Scheurwegs et al., , 2016), managing chronic diseases (Sheikhalishahi et al., , 2019), and pharmacovigilance (Stang et al., , 2010).
-While these different modalities serve as complementary data sources, there is significant overlap and correlation among these data (Qiao et al.,, 2019).
+While these different modalities serve as complementary data sources, there is significant overlap and correlation among these modalities (Qiao et al., , 2019).
 Joint representation of both structured and narrative data into a more manageable low-dimensional space where similar features are grouped closely can significantly improve the utility of both data types.
 This representation learning technique has gained popularity for its ability to capture and represent the intricate relationships among various EHR features.
-Despite the extensive research on EHR feature representation, most existing studies have primarily focused on either structured (Choi et al., 2016a, ; Kartchner et al.,, 2017; Hong et al.,, 2021; Zhou et al.,, 2022) or unstructured data modalities (De Vine et al.,, 2014; Choi et al., 2016b, ; Beam et al.,, 2019; Alsentzer et al.,, 2019; Huang et al.,, 2020; Lehman and Johnson,, 2023).
-For instance, Alsentzer et al., (2019) adapted the BERT model (Devlin et al.,, 2019) to the clinical domain by training on the MIMIC-III clinical notes (Johnson et al.,, 2016).
-It was extremely time/resource-consuming to train such a black-box model.
-De Vine et al., (2014) aligned free-text features with the Unified Medical Language System (UMLS)’s common concept unique identifier (CUI) space (McInnes et al.,, 2007).
-They then utilized the skip-gram algorithm (Mikolov et al.,, 2013) on concatenated concept documents to learn the CUI embeddings.
-Similarly, Choi et al., 2016b and Beam et al., (2019) employed singular value decomposition (SVD) on a pointwise mutual information (PMI) matrix derived from CUI co-occurrences in unstructured text to generate CUI embeddings.
-This approach was also adopted by Hong et al., (2021) for code embeddings.
-This SVD-PMI algorithm not only ensures scalability but also upholds data privacy through the use of aggregate co-occurrence data, offering a scalable variant of the skip-gram algorithm (Levy and Goldberg,, 2014).
-Its interpretability is further highlighted in the dynamic log-linear topic model developed by Arora et al., (2016).
-Recent studies have emerged on leveraging multimodal EHR features for enhanced predictive modeling, as demonstrated by the work of Khadanga et al., (2019); Zhang et al., (2020); Bardak and Tan, (2021); Gan et al., (2023).
-Specifically, Gan et al., (2023) enhanced code and CUI embeddings by employing the SVD-PMI algorithm, facilitating the integration of structured and unstructured data.
-Despite these advancements, methods relying on straightforward data merging may not fully account for the intricate interactions within multimodal data, potentially leading to biases.
-This issue will be further examined in our theoretical (Section 3) and numerical analyses (Sections 4 and 5).
-Addressing these limitations, Liu et al., (2022) introduced a multimodal pre-trained language model incorporating a cross-attention mechanism to enrich EHR representations across structured and unstructured data domains.
-Additionally, the domain has seen progress in multimodal contrastive learning strategies, as evidenced by research from Li and Gao, (2022); Yin et al., (2023); Wang et al., (2023).
-These strategies, drawing inspiration from successful vision-language models like the Contrastive Language-Image Pre-Training (CLIP) by Radford et al., (2021), aim to train unified representations of diverse data modalities.
-However, these approaches, grounded in deep neural networks, face challenges due to their “black-box” nature, including a lack of theoretical underpinning, computational complexity, and concerns over privacy since they necessitate access to individual patient data.
-These factors contribute to their limited applicability in the context of EHR data.
-While some theoretical analyses of multimodal learning exist, their applicability to EHR data has been limited.
+Despite extensive research on EHR feature representation, most studies have focused on datasets with a single modality, including structured (Choi et al., 2016a, ; Kartchner et al., , 2017; Hong et al., , 2021; Zhou et al., , 2022) and unstructured data (De Vine et al., , 2014; Choi et al., 2016b, ; Beam et al., , 2019; Alsentzer et al., , 2019; Huang et al., , 2020; Lehman and Johnson, , 2023).
+For example, Alsentzer et al., (2019) adapted BERT (Devlin et al., , 2019) to the clinical domain by pretraining on MIMIC-III notes (Johnson et al., , 2016), a process that was highly resource-intensive.
+De Vine et al., (2014) aligned free-text with the UMLS Concept Unique Identifier (CUI) space (McInnes et al., , 2007) and applied skip-gram (Mikolov et al., , 2013) to learn embeddings, while Choi et al., 2016b , Beam et al., (2019); Hong et al., (2021); Levy and Goldberg, (2014); Arora et al., (2016) used singular value decomposition (SVD) on pointwise mutual information (PMI) matrices.
+Building on these unimodal advances, recent work has explored multimodal EHR representations that integrate both structured and unstructured information (Khadanga et al., , 2019; Zhang et al., , 2020; Bardak and Tan, , 2021; Gan et al., , 2025).
+Joint representation is preferable to treating all data as a single modality because, although structured codes and unstructured text encode overlapping clinical concepts, they possess fundamentally different statistical forms.
+Ignoring these cross-modality differences can obscure modality-specific information, discard complementary signal, and amplify systematic bias (see further discussion in Section 3.2).
+By contrast, multimodal integration can exploit the complementary strengths of distinct modalities, yielding representations that are not only richer but also less biased than those obtained from unimodal analyses.
+For instance, information that is noisy or missing in one modality may be corroborated by the other, thereby stabilizing estimation and enhancing generalization.
+In this paper, we develop a theoretical framework (Section 3) and provide empirical evidence (Sections 4 and 5) to demonstrate that multimodal analysis addresses these challenges and produces more reliable and interpretable representations of EHR data.
+To address these limitations, Liu et al., (2022) proposed a multimodal pre-trained language model that incorporates a cross-attention mechanism to enrich EHR representations across structured and unstructured domains.
+More recently, multimodal contrastive learning strategies have also been explored (Li and Gao, , 2022; Yin et al., , 2023; Wang et al., , 2023).
+Inspired by the success of vision–language models such as CLIP (Radford et al., , 2021), these approaches aim to learn unified representations of heterogeneous data modalities.
+From a statistical perspective, contrastive learning can be interpreted through the lens of noise-contrastive estimation (Gutmann and Hyvärinen, , 2010), in which a parametric score function distinguishes positive’ pairs drawn from the joint distribution from negative’ pairs drawn from the product of marginals.
+The popular InfoNCE loss (Oord et al., , 2018) maximizes a lower bound on mutual information and is closely related to low-rank approximations of the PMI matrix (Levy and Goldberg, , 2014).
+Despite their promise, these methods remain limited by their reliance on deep neural networks, which are often criticized for their “black-box” nature.
+Practical challenges include the absence of rigorous theoretical guarantees, high computational cost, and privacy concerns due to the need for patient-level data.
+Together, these issues constrain the applicability of existing multimodal contrastive learning frameworks in the context of EHR data.
+While some theoretical analyses of multimodal learning exist, their applicability to sparse discrete EHR data has been limited.
 Groundbreaking studies like Huang et al., (2021) have illustrated the benefits of multimodal learning, showing that learning across multiple modalities can reduce population risk compared to single-modality methods.
@@
 However, these studies do not directly apply to the unique discrete feature structure of EHR data, leaving an unaddressed theoretical gap in understanding multimodal contrastive learning’s application in healthcare.
-Bridging this gap is vital, as it lays a solid foundation for multimodal contrastive learning’s development and implementation in healthcare, maximizing its potential to improve patient care and drive medical research forward.
+Bridging this gap is vital, as it lays a foundation for multimodal contrastive learning’s development and implementation in healthcare, maximizing its potential to improve care and drive medical research forward.
+Complementary lines of research have attempted to address related challenges from different angles.
+For instance, exponential family principal component analysis (PCA) (Collins et al., , 2001) and nonlinear PCA (Hinton and Salakhutdinov, , 2006) extend classical PCA to non-Gaussian and nonlinear settings, while kernel PCA (KPCA) (Gupta et al., , 2019) provides additional flexibility.
+Yet, handling multimodal, sparse, and privacy-sensitive data remains an open challenge.
+In parallel, the statistical literature on integrative dimension reduction, such as JIVE (Lock et al., , 2013), AJIVE (Feng et al., , 2018), and other multi-view factor models (Zhou et al., , 2015; Yang and Michailidis, , 2016; Yi et al., , 2023, e.g.,), provides tools for modeling shared and modality-specific structures.
+Our framework shares this goal but is tailored to discrete EHR features and prioritizes privacy-preserving learning via co-occurrence data.
 To overcome these limitations, we introduce the Contrastive Learning Algorithm for Integrated Multimodal Electronic health records (CLAIME).
-Our findings confirm that CLAIME is not only an effective tool for deriving multimodal EHR feature representations but also a method that respects privacy by requiring only aggregated data.
+Our findings confirm that CLAIME is not only effective for deriving multimodal EHR feature representations but also respects privacy by requiring only aggregated data.
 Additionally, we propose a novel multimodal feature embedding generative model (2.1) in Section 2.1, designed to enhance statistical analysis of multimodal EHR data.
 This model is notable for its interpretability and accurate portrayal of EHR data generation.
-It distinguishes itself from earlier word vector generative models (Arora et al.,, 2016, 2018; Lu et al.,, 2023; Xu et al.,, 2022) by (1) enabling the integration of multimodal EHR features, and (2) allowing patient heterogeneity by incorporating error terms specific to patients, thus increasing the model’s robustness.
+It distinguishes itself from earlier word vector generative models (Arora et al., , 2016, 2018; Lu et al., , 2023; Xu et al., , 2023) by (1) enabling the integration of multimodal EHR features, and (2) allowing patient heterogeneity by incorporating error terms specific to patients, thus increasing the model’s robustness.
 Within this generative framework, we validate the consistency of the CLAIME algorithm and clarify the relationship between multimodal feature embeddings and a multimodal PMI matrix.
@@
 Our research also fills a theoretical void in the analysis of multimodal contrastive learning for EHR data.
-The rest of the paper is structured as follows.
-Section 2 elaborates on the proposed method.
-Section 3 presents the theoretical properties of our algorithm.
-In Section 4, simulation studies demonstrate the algorithm’s effectiveness under various configurations.
-Section 5 further validates our algorithm’s clinical applicability in EHR studies.
-Finally, Section 6 includes discussions.
+The rest of the paper is organized as follows.
+Section 2 introduces the proposed method, and Section 3 presents its theoretical properties.
+Sections 4 and 5 provide simulation and real-world evaluations, respectively.
+Section 6 concludes with a discussion.
 2 Method
-2.1 Notation
-For any matrix $\mathbf{A}$, let $\|\mathbf{A}\|$, $\|\mathbf{A}\|_{\max}$ and $\|\mathbf{A}\|_{\mathrm{F}}$ be its operator norm, entrywise maximum norm and Frobenius norm, respectively.
-We define $\mathbf{P}_{p}(\mathbf{A})$ as the top-$p$ right singular vectors of $\mathbf{A}$.
-When the right singular vectors are not unique, we choose arbitrary singular vectors.
-Denote $s_{j}(\mathbf{A})$ as the $j$-th largest singular value of $A$.
-Let $\mathcal{O}_{d,p}$ ($d\geq p$) denote a set of $d\times p$ orthonormal matrices.
-For two sequences of positive numbers $\{a_{k}\}$ and $\{b_{k}\}$, we write $a_{k}\lesssim b_{k}$ or $a_{k}=O(b_{k})$ or $b_{k}\gtrsim a_{k}$ or $b_{k}=\Omega(a_{k})$ if there exists some constant $C>0$ such that $a_{k}\leq Cb_{k}$ for all $k$.
-We denote $a_{k}\ll b_{k}$ or $a_{k}=o(b_{k})$ if $\lim_{k\rightarrow\infty}a_{k}/b_{k}=0$.
-For any positive integer $I$, let $[I]=\{1,2,\cdots,I\}$.
-We write $a\vee b$ and $a\wedge b$ to denote $\max(a,b)$ and $\min(a,b)$, respectively.
-We denote by $\bm{e}_{j}$ the $j$-th unit vector in $\mathbb{R}^{d}$ where its $j$-th element is one, and all other elements are zero.
+We begin by defining the notation.
+For any matrix $\mathbf{A}$, let $\|\mathbf{A}\|$, $\|\mathbf{A}\|_{\max}$, and $\|\mathbf{A}\|_{\mathrm{F}}$ denote its operator norm, entrywise maximum norm, and Frobenius norm, respectively.
+Let $\mathbf{P}_{p}(\mathbf{A})$ be the top-$p$ right singular vectors of $\mathbf{A}$, chosen arbitrarily if non-unique, and let $s_{j}(\mathbf{A})$ be its $j$-th largest singular value.
+Define $\mathcal{O}_{d,p}$ ($d\geq p$) as the set of $d\times p$ orthonormal matrices.
+For sequences $\{a_{n}\},\{b_{n}\}>0$, write $a_{n}\lesssim b_{n}$, $a_{n}=O(b_{n})$, $b_{n}\gtrsim a_{n}$, or $b_{n}=\Omega(a_{n})$ if $a_{n}\leq Cb_{n}$ for some $C>0$ and all $n$; write $a_{n}\ll b_{n}$ or $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\to 0$.
+Let $I]=\{1,2,\dots,I\}$ for any positive integer $I$, and write $a\vee b$ and $a\wedge b$ for $\max(a,b)$ and $\min(a,b)$, respectively.
+Denote by $\bm{e}_{j}$ the $j$-th unit vector in $\mathbb{R}^{d}$.
 2.1 Model Assumptions
-Assume that the collections of codes from the structured modality and CUIs from the unstructured modality are denoted by ${\cal W}^{(1)}:=[d_{1}]$ and ${\cal W}^{(2)}:=[d]\setminus[d_{1}]$ respectively, where $d=d_{1}+d_{2}$, with $d$ representing the total number of unique features across both modalities.
-Suppose that we have $n$ independent patients.
-For each patient $i$, the observed codes and CUIs are denoted as $\{w^{(1)}_{i,t}\}_{t\in[T_{i}^{(1)}]}$ and $\{w^{(2)}_{i,t}\}_{t\in[T_{i}^{(2)}]}$, where $w^{(1)}_{i,t}\in{\cal W}^{(1)}$ and $w^{(2)}_{i,t}\in{\cal W}^{(2)}$.
-The sizes of these sets are given by $T_{i}^{(1)}\geq 2$ and $T_{i}^{(2)}\geq 2$, respectively.
-We model the probability of observing specific codes and CUIs for the $i$-th patient based on their embeddings $\mathbf{v}_{w}^{\star}\in\mathbb{R}^{p}$ as follows:
-$\mathbb{P}(w^{(1)}_{i,t}=w|\bm{c}_{i},{\bm{\epsilon}}_{i}^{(1)})$ $=\frac{{\exp}(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w})}{\sum_{w^{\prime}\in{\cal W}^{(1)}}{\exp}(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}})},\ \ w\in{\cal W}^{(1)},t\in[T_{i}^{(1)}],\text{ and }$ (2.1)
-$\mathbb{P}(w^{(2)}_{i,t}=w|\bm{c}_{i},{\bm{\epsilon}}_{i}^{(2)})$ $=\frac{{\exp}(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w})}{\sum_{w^{\prime}\in{\cal W}^{(2)}}{\exp}(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}})},\ \ w\in{\cal W}^{(2)},t\in[T_{i}^{(2)}]\,.$
-Here $\bm{c}_{i}\sim N(\bm{0},\mathbf{I}_{p})$ represents a latent vector specific to patient $i$, reflecting their clinical state.
-The error terms, ${\bm{\epsilon}}^{(1)}_{i}=(\epsilon_{i,1},\dots,\epsilon_{i,d_{1}})^{\top}\sim N(\bm{0},{\bm{\Sigma}}_{1})$ and ${\bm{\epsilon}}^{(2)}_{i}=(\epsilon_{i,d_{1}+1},\dots,\epsilon_{i,d})^{\top}\sim N(\bm{0},{\bm{\Sigma}}_{2})$, account for patient-specific variations and address the variability not captured by the term $\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle$, where ${\bm{\Sigma}}_{1}$ and ${\bm{\Sigma}}_{2}$ are some unknown positive semi-definite matrices.
-2.3 CLAIME Algorithm
-We define code and CUI embedding matrices as $\mathbf{V}_{1}^{\star}=(\mathbf{v}_{1}^{\star},\ldots,\mathbf{v}_{d_{1}}^{\star})^{\top}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}^{\star}=(\mathbf{v}_{d_{1}+1}^{\star},\ldots,\mathbf{v}_{d}^{\star})^{\top}\in\mathbb{R}^{d_{2}\times p}$, respectively, and aim to infer $\mathbf{V}^{\star}=\big{(}\mathbf{V}_{1}^{\star\top},\mathbf{V}_{2}^{\star\top}\big{)}^{\top}\in\mathbb{R}^{d\times p}$.
-The embeddings should reflect clinical semantics, meaning that highly similar (e.g.
-rheumatoid arthritis and juvenile rheumatoid arthritis) or related (e.g.
-fasting glucose and type II diabetes) EHR entities should have close embeddings.
+[…]
+Figure 1: Illustration of multimodal EHR generation: A shared latent state $\bm{c}_{i}$ generates both structured codes (top) and unstructured CUIs (bottom), motivating (2.1).
+Our approach is motivated by real-world EHRs, where structured codes (e.g., diagnoses, medications) and unstructured concepts extracted from clinical notes (e.g., CUIs) offer complementary views of a patient’s health (Figure 1).
+During a hospital visit, patient $i$ generates two types of data: structured codes and CUIs, which are typically collected by different subsystems and subject to distinct noise processes.
+For example, structured codes often reflect billing practices, while CUIs are derived from free-text notes via Named Entity Recognition (NER) tools.
+This data processing pipeline aligns with standard practice for mapping text to UMLS concepts (Choi et al., 2016b, ; Beam et al., , 2019).
+We primarily focus on the two-modality case but also discuss extensions to more general multi-modality settings involving three or more modalities below.
+Accordingly, for each patient $i$, we observe two sets of features: $\{w^{(1)}_{i,t}\}_{t=1}^{T_{i}^{(1)}}$ for structured codes and $\{w^{(2)}_{i,t}\}_{t=1}^{T_{i}^{(2)}}$ for CUIs, where $w^{(1)}_{i,t}\in{\cal W}^{(1)}$ and $w^{(2)}_{i,t}\in{\cal W}^{(2)}$.
+Let ${\cal W}^{(1)}:=[d_{1}]$ and ${\cal W}^{(2)}:=[d]\setminus[d_{1}]$, where $d=d_{1}+d_{2}$ denotes the total number of distinct features across both modalities.
+To model the generative process of these features, we adopt the following latent variable model, which captures the shared structure between modalities while allowing for modality-specific noise:
+$\mathbb{P}(w^{(1)}_{i,t}=w\mid\bm{c}_{i},{\bm{\epsilon}}_{i}^{(1)})$ $=\frac{\exp(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w}^{(1)})}{\sum_{w^{\prime}\in{\cal W}^{(1)}}\exp(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}}^{(1)})},\quad w\in{\cal W}^{(1)},\ t\in[T_{i}^{(1)}],$ (2.1)
+$\mathbb{P}(w^{(2)}_{i,t}=w\mid\bm{c}_{i},{\bm{\epsilon}}_{i}^{(2)})$ $=\frac{\exp(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w}^{(2)})}{\sum_{w^{\prime}\in{\cal W}^{(2)}}\exp(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}}^{(2)})},\quad w\in{\cal W}^{(2)},\ t\in[T_{i}^{(2)}].$
+Here, $\bm{c}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{p})$ represents the latent clinical state of patient $i$, and $\mathbf{v}_{w}^{\star}\in\mathbb{R}^{p}$ is the ground-truth embedding of feature $w$.
+The modality-specific noise vectors ${\bm{\epsilon}}^{(M)}_{i}=(\epsilon_{i,w}^{(M)})_{w\in{\cal W}^{(M)}}\sim\mathcal{N}(\mathbf{0},{\bm{\Sigma}}_{M})$, for $M\in\{1,2\}$, capture variability not explained by the latent state, such as documentation inconsistencies or NLP extraction errors.
+This formulation reflects our core assumption: while structured and unstructured features are conditionally generated from a shared latent representation $\bm{c}_{i}$, they are distorted independently by modality-specific noise.
+The model thus supports: (1) integration of heterogeneous EHR data sources, (2) robustness to differing noise structures across modalities, and (3) inference of informative embeddings from only summary-level statistics.
+Remark 2.1.
+The assumption $\bm{c}_{i}\sim N(\mathbf{0},\mathbf{I}_{p})$ is used to support theoretical analysis, not as a requirement for the CLAIME algorithm itself.
+Our method does not rely on estimating individual $\bm{c}_{i}$ vectors.
+Instead, the Gaussian assumption enables analytical tractability in characterizing the expected co-occurrence patterns of EHR features, which in turn justify the use of SVD-based estimators in later sections.
+Similar analytical approximations have been derived under alternative latent structures, such as unit sphere random walks (Arora et al., , 2016, 2017) or scaled-Gaussian priors (Xu et al., , 2023).
+Finally, due to the rotational invariance of our model, the identity covariance assumption on $\bm{c}_{i}$ can be relaxed without affecting the core methodology.
+2.2 The CLAIME Algorithm
+We define code and CUI embedding matrices as $\mathbf{V}_{1}^{\star}=(\mathbf{v}_{1}^{\star},\ldots,\mathbf{v}_{d_{1}}^{\star})^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}^{\star}=(\mathbf{v}_{d_{1}+1}^{\star},\ldots,\mathbf{v}_{d_{1}+d_{2}}^{\star})^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d_{2}\times p}$, respectively, and aim to infer $\mathbf{V}^{\star}=\big{(}\mathbf{V}_{1}^{\star{\scriptscriptstyle\sf T}},\mathbf{V}_{2}^{\star{\scriptscriptstyle\sf T}}\big{)}^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d\times p}$.
+The embeddings are intended to reflect clinical semantics, meaning that highly similar (e.g., rheumatoid arthritis and juvenile rheumatoid arthritis) or related (e.g., fasting glucose and type II diabetes) EHR entities should be embedded close to each other.
+To motivate our algorithm design, recall that model (2.1) assumes EHR features from both modalities are generated from a shared latent patient embedding $\bm{c}_{i}$.
+This implies that features frequently co-occurring within the same patient are likely conditionally dependent through $\bm{c}_{i}$, and should have similar embeddings, whereas features from different patients are marginally independent and their embeddings should remain unaligned.
 Before introducing our algorithm, we first define aggregate co-occurrence matrices $\mathbf{C}^{(M,M)}$ and $\mathbf{D}^{(M,M^{\prime})}$ for $M,M^{\prime}\in\{1,2\}$ across different modalities as:
 $\mathbf{C}^{(M,M)}(w,w^{\prime})=\sum_{i=1}^{n}\mathbf{C}^{(M,M)}_{i,i}(w,w^{\prime}),\ \ \mathbf{D}^{(M,M^{\prime})}(w,w^{\prime})=\sum_{i=1}^{n}\mathbf{D}^{(M,M^{\prime})}_{i,i}(w,w^{\prime})$
-where $\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M)}]:t\neq s,w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ and $\mathbf{D}^{(M,M^{\prime})}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M^{\prime})}]:w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ for $i,j\in[n]$ and $M\in\{1,2\}$.
-Further, we define the marginal co-occurrence of $w\in{\cal W}^{(M)}$ as:
+where $\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M)}]:t\neq s,w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ and $\mathbf{D}^{(M,M^{\prime})}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M^{\prime})}]:w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ for $i,j\in[n]$.
+Further, we define the marginal occurrence count of $w\in{\cal W}^{(M)}$ as:
 $\gamma_{w}^{(M)}=\mathbf{C}^{(M,M)}(w,\cdot)=\sum_{w^{\prime}\in{\cal W}^{(M)}}\mathbf{C}^{(M,M)}(w,w^{\prime}).$ (2.2)
-We introduce $S_{q}^{(M)}=(n^{-1}\sum_{i=1}^{n}(T_{i}^{(M)})^{q})^{1/q}$ for $q\geq 1$ and $S_{1}^{(1,2)}=n^{-1}\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}$ as scaling factors.
-CLAIME utilizes the multimodal contrastive learning loss defined as:
-$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})=\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}\sum_{i\neq j}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{j,s}}\rangle}{\gamma_{w_{i,t}}^{(1)}\gamma_{w_{j,s}}^{(2)}}$ (2.3)
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{1}{nS_{1}^{(1,2)}}\sum_{i=1}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{i}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{i,s}}\rangle}{\gamma_{w_{i,t}}^{(1)}\gamma_{w_{i,s}}^{(2)}}+\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\top}\|_{\mathrm{F}}^{2}.$
-Here $\mathbf{V}_{1}=(\mathbf{v}_{1},\ldots,\mathbf{v}_{d_{1}})^{\top}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}=(\mathbf{v}_{d_{1}+1},\ldots,\mathbf{v}_{d})^{\top}\in\mathbb{R}^{d_{2}\times p}$, $\lambda>0$ serves as a regularization coefficient, $\gamma_{w}^{(M)},M=1,2$ are weights chosen based on the frequency of $w$, as defined in (2.2).
-Our theoretical analysis in Section 3 motivates the choice of $\gamma_{w}^{(M)}$ to guide the minimizer towards $\mathbf{V}^{\star}$.
-The essence of the multimodal contrastive learning loss in CLAIME is to enhance the representation of similar features across different modalities by bringing them closer together while distancing those that are dissimilar.
-In the context of EHR data, this translates to aligning the embeddings of codes and CUIs that have clinical correlations and separating those that do not.
-The aim is to maximize the inner product of embeddings for features that co-occur within the same patient’s data.
-Remark 2.1.
-Our CLAIME framework can be naturally extended to non-lienar loss functions. Note that our CLAIME loss function (2.3) can be written as
-$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=-\frac{1}{\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}}\sum_{i=1}^{n}\quantity{T_{i}^{(1)}T_{i}^{(2)}s_{ii}-\alpha_{T}\sum_{j:j\neq i}^{n}T_{i}^{(1)}T_{j}^{(2)}s_{ij}}+R(\mathbf{V}_{1},\mathbf{V}_{2}),$
-where $R(\mathbf{V}_{1},\mathbf{V}_{2})$ is a smooth regularizer, and
-$\alpha_{T}=\frac{\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}}{\sum_{i\neq j}^{n}T_{i}^{(1)}T_{j}^{(2)}},\ \ s_{ij}=\frac{1}{T_{i}^{(1)}T_{j}^{(2)}}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{j,s}}\rangle}{\gamma_{w_{i,t}}^{(1)}\gamma_{w_{j,s}}^{(2)}},\ \ \ \ \ \ i,j\in[n].$
-One can consider the following non-linear loss function analogous to the CLIP loss function (Radford et al.,, 2021):
-$\mathcal{L}_{\textnormal{CLAIME}}^{\prime}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=-\frac{1}{\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}}\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}\log\frac{\exp(s_{ii}/\tau)}{\sum_{j:j\neq i}^{n}T_{i}^{(1)}T_{j}^{(2)}\exp(s_{ij}/\tau)}+R(\mathbf{V}_{1},\mathbf{V}_{2}),$
-which becomes equivalent to the loss (2.3) when $\tau\to\infty$.
-To obtain $\mathbf{\widehat{V}}=(\mathbf{\widehat{V}}_{1}^{\top},\mathbf{\widehat{V}}_{2}^{\top})^{\top}=\arg\min\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ efficiently in practice, we note that $\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ can be expressed in terms of pair-wise co-occurrences of concepts as:
-$\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}^{\star}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{j\neq i}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{j,s}=w^{\prime})$ (2.4)
-$\quad-\frac{1}{nS_{1}^{(1,2)}}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}^{\star}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{i}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{i,s}=w^{\prime})+\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\top}\|_{\mathrm{F}}^{2}\,.$
+To derive clinically meaningful embeddings that reflect modality-specific co-occurrence patterns, we propose a contrastive learning algorithm aligned with model (2.1).
+Contrastive learning encourages embeddings of positive pairs to be similar and those of negative pairs to be dissimilar.
+In our setting, code-CUI pairs from the same patient are treated as positive, while pairs from different patients as negative.
+This naturally aligns with our generative assumptions and enables learning from aggregated co-occurrence data.
+Specifically, we define the multimodal contrastive learning loss as
+$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})=\frac{\sum_{i,j:i\neq j}s_{ij}}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}-\frac{\sum_{i=1}^{n}s_{ii}}{nS_{2}^{(1,2)2}}+\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\|_{\mathrm{F}}^{2},$ (2.3)
+where $s_{ij}=\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{j,s}}\rangle}{\gamma_{w_{i,t}^{(1)}}^{(1)}\gamma_{w_{j,s}^{(2)}}^{(2)}}$.
+Here, $\mathbf{V}_{1}=(\mathbf{v}_{1},\ldots,\mathbf{v}_{d_{1}})^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}=(\mathbf{v}_{d_{1}+1},\ldots,\mathbf{v}_{d_{1}+d_{2}})^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d_{2}\times p}$, $\lambda>0$ serves as a regularization coefficient, $\gamma_{w}^{(M)},M=1,2$ are weights chosen based on the frequency of $w$, as defined in (2.2).
+The quantities $S_{q}^{(M)}=(n^{-1}\sum_{i=1}^{n}(T_{i}^{(M)})^{q})^{1/q}$ and $S_{q}^{(1,2)}=(n^{-1}\sum_{i\in[n]}(T_{i}^{(1)})^{q/2}(T_{i}^{(2)})^{q/2})^{1/q}$ for $q\geq 1$ are empirical moments and cross-moments of $T_{i}^{(M)}$, and correspond to the number of terms in the two summation components of the loss.
+Our analysis in Section 3 motivates the choice of $\gamma_{w}^{(M)}$ to guide the minimizer towards $\mathbf{V}^{\star}$.
+This loss encourages code-CUI pairs that co-occur within the same patient (second term) to be closely embedded, while penalizing similarity between pairs from different patients (the first term).
+The regularization term $\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\|_{\mathrm{F}}^{2}$ ensures numerical stability and controls the embedding magnitude.
+By optimizing this loss, the model learns to align embeddings based on co-occurrence patterns, which signal shared clinical relevance.
+To efficiently obtain $\mathbf{\widehat{V}}=(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T},\mathbf{\widehat{V}}_{2}^{\scriptscriptstyle\sf T})^{\scriptscriptstyle\sf T}=\arg\min\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$, we define normalized count vectors for each patient $i$ and modality $M\in\{1,2\}$: $\mathbf{w}_{i}^{(M)}\in\mathbb{R}^{d_{M}}$ with $\mathbf{w}_{i,w}^{(M)}=\frac{1}{\gamma_{w}^{(M)}}\sum_{t\in[T_{i}^{(M)}]}\mathbb{I}(w_{i,t}^{(M)}=w)$.
+The patient-level representation for each modality becomes $\mathbf{V}_{1}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(1)}$ and $\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(2)}\in\mathbb{R}^{p}$.
+Using this, the loss (2.3) (excluding regularization) can be expressed by pair-wise co-occurrences of concepts as:
+$\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{1}+d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{j,s}=w^{\prime})$
+$\quad-\frac{1}{nS_{2}^{(1,2)2}}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{1}+d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{i}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{i,s}=w^{\prime})$
+$=\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}(\mathbf{V}_{1}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(1)})^{\scriptscriptstyle\sf T}(\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\mathbf{w}_{j}^{(2)})-\frac{1}{nS_{2}^{(1,2)2}}\sum_{i=1}^{n}(\mathbf{V}_{1}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(1)})^{\scriptscriptstyle\sf T}(\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(2)}).$
+The first term captures the average similarity of negative (cross-patient) pairs, while the second reflects positive (same-patient) pairs.
+Minimizing this loss encourages within-patient alignment while discouraging cross-patient alignment.
 Subsequently, via arguments given in Supplementary S2, we have the following proposition.
@@
 We have
-$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=\frac{\lambda}{2}\Big{\|}\mathbf{V}_{1}\mathbf{V}_{2}^{\top}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant}),$
+$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=\frac{\lambda}{2}\Big{\|}\mathbf{V}_{1}\mathbf{V}_{2}^{\scriptscriptstyle\sf T}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant}),$
 where $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}=\{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})\}_{w\in{\cal W}^{(1)},w^{\prime}\in{\cal W}^{(2)}}$ with
-$\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ $:=\frac{\mathbf{C}^{(1,1)}(\cdot,\cdot)\mathbf{C}^{(2,2)}(\cdot,\cdot)}{\mathbf{C}^{(1,1)}(w,\cdot)\mathbf{C}^{(2,2)}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{D}^{(1,2)}(w,w^{\prime})}{\mathbf{D}^{(1,2)}(\cdot,\cdot)}-\frac{\mathbf{C}^{(c)}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}),$
-and $\mathbf{C}^{(c)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j\neq i}^{n}\mathbf{D}^{(1,2)}_{i,j}(w,w^{\prime}).$
-Proposition 2.1 shows that $\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ can be related to the SVD of an empirical association matrix, where its element $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ estimates the association between the features $w$ and $w^{\prime}$.
-We will later demonstrate the convergence of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ to the population PMI matrix in Section 3.
-As a result, $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ can be viewed as a modified estimator of the population PMI matrix.
-The final embeddings $\mathbf{\widehat{V}}_{1}\mathbf{\widehat{V}}_{2}^{\top}$ are inferred through a rank-$p$ SVD of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, preserving data privacy and offering a scalable estimation.
-To be more specific, let $\widehat{\mathbf{U}}_{1}\widehat{\bm{\Lambda}}\widehat{\mathbf{U}}_{2}$ denote the rank-$p$ SVD of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, where $\widehat{\mathbf{U}}_{1}\in\mathbb{R}^{d_{1}\times p}$ and $\widehat{\mathbf{U}}_{2}\in\mathbb{R}^{d_{2}\times p}$ are the matrices of left and right singular vectors, respectively, and $\widehat{\bm{\Lambda}}\in\mathbb{R}^{p\times p}$ is a diagonal matrix with its diagonal elements being the top $p$ singular values.
+$\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime}):=\frac{\mathbf{C}^{(1,1)}(\cdot,\cdot)\mathbf{C}^{(2,2)}(\cdot,\cdot)}{\mathbf{C}^{(1,1)}(w,\cdot)\mathbf{C}^{(2,2)}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{D}^{(1,2)}(w,w^{\prime})}{\mathbf{D}^{(1,2)}(\cdot,\cdot)}-\frac{\mathbf{C}^{(c)}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}),$
+and $\mathbf{C}^{(c)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}\mathbf{D}^{(1,2)}_{i,j}(w,w^{\prime}).$
+Proposition 2.1 shows that $\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ reduces to a rank-constrained approximation of an empirical association matrix $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, where each entry $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ quantifies the association between features $w$ and $w^{\prime}$.
+Section 3 establishes that this matrix converges to the population-level PMI matrix under mild assumptions.
+To estimate the embeddings, we compute the rank-$p$ SVD of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, denoted as $\widehat{\mathbf{U}}_{1}\widehat{\bm{\Lambda}}\widehat{\mathbf{U}}_{2}^{{\scriptscriptstyle\sf T}}$, where $\widehat{\mathbf{U}}_{1}\in\mathbb{R}^{d_{1}\times p}$ and $\widehat{\mathbf{U}}_{2}\in\mathbb{R}^{d_{2}\times p}$ are the matrices of left and right singular vectors, respectively, and $\widehat{\bm{\Lambda}}\in\mathbb{R}^{p\times p}$ is a diagonal matrix with its diagonal elements being the top $p$ singular values.
 Then, we set $\mathbf{\widehat{V}}_{1}=\widehat{\mathbf{U}}_{1}\widehat{\bm{\Lambda}}^{1/2}$ and $\mathbf{\widehat{V}}_{2}=\widehat{\mathbf{U}}_{2}\widehat{\bm{\Lambda}}^{1/2}$.
-It is worth noting that the selection of the regularization parameter $\lambda$ does not play a crucial role, and for the sake of simplicity, we will assign it a value of $1$ in our upcoming numerical analyses.
-2.4 Comparison between CLAIME and Simple Concatenation
-We next contrast CLAIME with the simple approach of ignoring between-modality differences between different modalities.
+From Proposition 2.1, the scalar $\lambda$ in (2.3) controls the scale of the embeddings via the magnitude of singular values, but not their subspace.
+Since downstream similarity metrics like cosine similarity are scale-invariant, we may fix $\lambda=1$ for simplicity.
+Our formulation is inspired by contrastive learning, which promotes high similarity among positive pairs and low similarity among negatives.
+Unlike vision-language models such as CLIP (Radford et al., , 2021), it operates on discrete, high-dimensional EHR features rather than continuous image or text embeddings, and relies only on aggregated co-occurrence statistics instead of raw patient-level data.
+Remark 2.2 (Nonlinear extension).
+The CLAIME framework can be naturally extended to a softmax-based contrastive loss similar to that used in CLIP:
+$\mathcal{L}_{\textnormal{CLAIME}}^{\prime}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=-\frac{1}{n(S_{2}^{(1,2)})^{2}}\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}\log\frac{\exp(s_{ii}/(T_{i}^{(1)}T_{i}^{(2)}\eta))}{\sum_{j:j\neq i}T_{i}^{(1)}T_{j}^{(2)}\exp(\alpha_{i}s_{ij}/(T_{i}^{(1)}T_{j}^{(2)}\eta))}+R^{\prime}(\mathbf{V}_{1},\mathbf{V}_{2}),$ (2.4)
+where $R^{\prime}$ is a smooth regularizer, $\eta>0$ the temperature, and $\alpha_{i}$ depends on $(T_{i}^{(M)})_{i\in[n],M=1,2}$.
+In the high-temperature limit ($\eta\to\infty$), the loss reduces to the linear CLAIME loss (2.3), up to a constant and regularization.
+Details are provided in Supplementary S2.2.
+Remark 2.3 (Extending CLAIME to more than two modalities).
+EHR data often consist of multiple heterogeneous streams.
+For instance, codified data alone encompass diagnoses, procedures, medications, and laboratory results, each of which can be treated as a distinct modality.
+Similarly, NLP-derived CUIs can be partitioned by semantic types such as Disease or Syndrome and Diagnostic Procedure.
+To accommodate such settings, we generalize CLAIME beyond two modalities in Supplementary S6 by estimating cross-modal PMI matrices for every pair of modalities and jointly learning embeddings over all sources.
+Further extensions to continuous-valued sources such as genomics, imaging, or clinical tests will require suitable generative models for continuous data.
+2.3 Comparison between CLAIME and Baseline Approaches
+We next contrast CLAIME with the baseline approaches of ignoring between-modality differences.
 Dealing with multimodal data often presents difficulties, leading to conventional methods that overlook the differences between various modalities.
@@
 However, this rudimentary treatment of multimodal data may lead to substantial bias due to the inherent heterogeneity between different modalities.
-To illustrate, consider the concatenated data for the $i$-th patient represented as
-$\{w_{i,t}\}_{t\in[T_{i}]}=\Big{(}w_{i,1}^{(1)},\dots,w_{i,T_{i}^{(1)}}^{(1)},w_{i,1}^{(2)},\dots,w_{i,T_{i}^{(2)}}^{(2)}\Big{)},\text{ where }T_{i}=T_{i}^{(1)}+T_{i}^{(2)}\,.$
-A popular method to handle such data is the SVD-PMI algorithm, as referenced in (Levy and Goldberg,, 2014; Gan et al.,, 2023).
+To illustrate, consider the concatenated data for the $i$-th patient represented as $\{w_{i,t}\}_{t\in[T_{i}]}=\Big{(}w_{i,1}^{(1)},\dots,w_{i,T_{i}^{(1)}}^{(1)},w_{i,1}^{(2)},\dots,w_{i,T_{i}^{(2)}}^{(2)}\Big{)},\text{ where }T_{i}=T_{i}^{(1)}+T_{i}^{(2)}\,.$ A popular method to handle such data is the SVD-PMI algorithm, as referenced in Levy and Goldberg, (2014); Gan et al., (2025).
 In this context, we establish the co-occurrence matrices for the concatenated dataset as follows:
-$\mathbf{C}=\begin{bmatrix}\mathbf{C}^{(1,1)}&\mathbf{D}^{(1,2)}\\ \mathbf{D}^{(2,1)}&\mathbf{C}^{(2,2)}\end{bmatrix},$
-where $\mathbf{C}^{(M,M)}$ and $\mathbf{D}^{(M,M^{\prime})}$ are defined in Section 2.3.
-Subsequently, the empirical concatenated PMI matrix, denoted as $\widehat{\mathbb{PMI}}=\{\widehat{\mathbb{PMI}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$, is formulated as
-$\widehat{\mathbb{PMI}}(w,w^{\prime})$ $=\log\frac{\mathbf{C}(w,w^{\prime})\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)},$ (2.5)
-where $\mathbf{C}(w,\cdot)=\sum_{w^{\prime}=1}^{d}\mathbf{C}(w,w^{\prime})$ and $\mathbf{C}(\cdot,\cdot)=\sum_{w=1}^{d}\mathbf{C}(w,\cdot)$.
-Following this, we conduct a rank-$p$ eigen-decomposition of $\widehat{\mathbb{PMI}}$, represented as $\mathbf{\widehat{U}}_{\textnormal{Con}}\widehat{\bm{\Lambda}}_{\textnormal{Con}}\mathbf{\widehat{U}}_{\textnormal{Con}}$.
+$\mathbf{C}=\begin{bmatrix}\mathbf{C}^{(1,1)}&\mathbf{D}^{(1,2)}\\ \mathbf{D}^{(2,1)}&\mathbf{C}^{(2,2)}\end{bmatrix}$.
+Subsequently, the empirical concatenated PMI matrix, denoted as $\widehat{\mathbb{PMI}}=\{\widehat{\mathbb{PMI}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$, is formulated as $\widehat{\mathbb{PMI}}(w,w^{\prime})=\log\frac{\mathbf{C}(w,w^{\prime})\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)}$, where $\mathbf{C}(w,\cdot)=\sum_{w^{\prime}=1}^{d}\mathbf{C}(w,w^{\prime})$ and $\mathbf{C}(\cdot,\cdot)=\sum_{w=1}^{d}\mathbf{C}(w,\cdot)$.
+Following this, we conduct a rank-$p$ eigen-decomposition of $\widehat{\mathbb{PMI}}$, represented as $\mathbf{\widehat{U}}_{\textnormal{Con}}\widehat{\bm{\Lambda}}_{\textnormal{Con}}\mathbf{\widehat{U}}_{\textnormal{Con}}^{{\scriptscriptstyle\sf T}}$.
 The estimator of $\mathbf{V}^{\star}$ is then achieved by setting $\mathbf{\widehat{V}}_{\textnormal{Con}}=\mathbf{\widehat{U}}_{\textnormal{Con}}\widehat{\bm{\Lambda}}_{\textnormal{Con}}^{1/2}$.
@@
 $\mathcal{L}_{\textnormal{CL}}(\mathbf{V})=-\frac{1}{\sum_{i=1}^{n}T_{i}(T_{i}-1)}\sum_{i=1}^{n}\sum_{t\in[T_{i}]}\sum_{s\in[T_{i}]\setminus\{t\}}\frac{\langle\mathbf{v}_{w_{i,t}},\mathbf{v}_{w_{i,s}}\rangle}{\gamma_{w_{i,t}}\gamma_{w_{i,s}}}$ (2.5)
-$\quad\quad+\frac{1}{\sum_{i=1}^{n}\sum_{j\neq i}^{n}T_{i}T_{j}}\sum_{i\neq j}\sum_{t\in[T_{i}]}\sum_{s\in[T_{j}]}\frac{\langle\mathbf{v}_{w_{i,t}},\mathbf{v}_{w_{j,s}}\rangle}{\gamma_{w_{i,t}}\gamma_{w_{j,s}}}+\frac{\lambda}{2}\|\mathbf{V}\mathbf{V}^{\top}\|_{\mathrm{F}}^{2},$
+$\quad\quad+\frac{1}{\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}T_{i}T_{j}}\sum_{i,j:i\neq j}\sum_{t\in[T_{i}]}\sum_{s\in[T_{j}]}\frac{\langle\mathbf{v}_{w_{i,t}},\mathbf{v}_{w_{j,s}}\rangle}{\gamma_{w_{i,t}}\gamma_{w_{j,s}}}+\frac{\lambda}{2}\|\mathbf{V}\mathbf{V}^{\scriptscriptstyle\sf T}\|_{\mathrm{F}}^{2},$
 where $\gamma_{w}=\mathbf{C}(w,\cdot)$.
@@
 Here, we define
-$\mathbf{D}=\begin{bmatrix}\mathbf{C}^{(1)}&\mathbf{C}^{(c)}\\ \mathbf{C}^{(c)\top}&\mathbf{C}^{(2)}\end{bmatrix}\text{ with }\mathbf{C}^{(M)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j\neq i}^{n}\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})\text{ for }w,w^{\prime}\in{\cal W}^{(M)}.$
+$\mathbf{D}=\begin{bmatrix}\mathbf{C}^{(1)}&\mathbf{C}^{(c)}\\ \mathbf{C}^{(c){\scriptscriptstyle\sf T}}&\mathbf{C}^{(2)}\end{bmatrix}\text{ with }\mathbf{C}^{(M)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})\text{ for }w,w^{\prime}\in{\cal W}^{(M)}.$
 Proposition 2.2.
-We have
-$\mathcal{L}_{\textnormal{CL}}(\mathbf{V})$ $=\frac{\lambda}{2}\Big{\|}\mathbf{V}\mathbf{V}^{\top}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CL}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant}),$
-where $\widehat{\mathbb{PMI}}_{\textnormal{CL}}=\{\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$ with
-$\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime}):=\frac{\mathbf{C}(\cdot,\cdot)\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{C}(w,w^{\prime})}{\mathbf{C}(\cdot,\cdot)}-\frac{\mathbf{D}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}).$
+We have $\mathcal{L}_{\textnormal{CL}}(\mathbf{V})=\frac{\lambda}{2}\Big{\|}\mathbf{V}\mathbf{V}^{\scriptscriptstyle\sf T}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CL}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant})$, where $\widehat{\mathbb{PMI}}_{\textnormal{CL}}=\{\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime})\}_{w,w^{\prime}\in[d]}$ with
+$\widehat{\mathbb{PMI}}_{\textnormal{CL}}(w,w^{\prime}):=\frac{\mathbf{C}(\cdot,\cdot)\mathbf{C}(\cdot,\cdot)}{\mathbf{C}(w,\cdot)\mathbf{C}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{C}(w,w^{\prime})}{\mathbf{C}(\cdot,\cdot)}-\frac{\mathbf{D}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}).$
 Proposition 2.2 reveals that obtaining $\mathbf{\widehat{V}}_{\textnormal{CL}}$ by minimizing the loss $\mathcal{L}_{\textnormal{CL}}$ is essentially equivalent to applying PCA on the matrix $\widehat{\mathbb{PMI}}_{\textnormal{CL}}$.
@@
 3 Theoretical Analysis
-In this section, we investigate the theoretical properties of CLAIME and draw comparisons with the Concate and CL methods.
-First, let us establish some fundamental concepts.
-For $w\in{\cal W}^{(M)}$, and $t\in[T_{i}^{(M)}]$, denote $X_{i,w}^{(M)}(t)=\mathbb{I}\{w^{(M)}_{i,t}=w\}$.
-According to the model (2.1), conditioned on $\bm{c}_{i}$ and ${\bm{\epsilon}}_{i}^{(M)}$, the variable $X_{i,w}^{(M)}(t)$ follows a Bernoulli distribution, with the probability of occurrence given by
-$p_{i,w}^{(M)}=\mathbb{E}[X_{i,w}^{(M)}(t)|\bm{c}_{i},{\bm{\epsilon}}_{i}^{(M)}]=\frac{\exp(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w})}{\sum_{w^{\prime}\in{\cal W}^{(M)}}\exp(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}})}\,.$ (3.1)
-Here, the probability $p_{i,w}^{(M)}$ is a function of the discourse vector $\bm{c}_{i}$ and the noise term ${\bm{\epsilon}}_{i}^{(M)}$, and is independent of $t$.
-Next, we denote the expected value of $p_{i,w}^{(M)}$ as $p_{w}^{(M)}=\mathbb{E}[p_{i,w}^{(M)}]$.
-Further, we define the co-occurrence of a feature $w\in{\cal W}^{(M)}$ at time $t$ for the $i$-th patient and feature $w^{\prime}\in{\cal W}^{(M^{\prime})}$ at time $s$ for the $j$-th patient as $X_{i,j,w,w^{\prime}}^{(M,M^{\prime})}(t,s)=\mathbb{I}\{w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}$ for $t\neq s$.
-This variable follows a Bernoulli distribution, with the probability of occurrence denoted as $p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}$.
-Furthermore, we define the expected value for the case of $i=j$ as $p_{w,w^{\prime}}^{(M,M^{\prime})}=\mathbb{E}[p_{i,i,w,w^{\prime}}^{(M,M^{\prime})}]$.
-It is important to note that when $i\neq j$, the expected value $\mathbb{E}[p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}]$ equals the product $p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}$.
-Now, we introduce the population PMI matrix $\mathbb{PMI}$ defined as
-$\mathbb{PMI}:=\begin{pmatrix}\mathbb{PMI}^{(1,1)}&\mathbb{PMI}^{(1,2)}\\ \mathbb{PMI}^{(2,1)}&\mathbb{PMI}^{(2,2)}\end{pmatrix}\text{ with }\mathbb{PMI}^{(M,M^{\prime})}(w,w^{\prime})=\log\frac{p_{w,w^{\prime}}^{(M,M^{\prime})}}{p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}}$ (3.4)
-for $w\in{\cal W}^{(M)}$, $w^{\prime}\in{\cal W}^{(M^{\prime})}$ with $M,M^{\prime}\in\{1,2\}$.
-Before proceeding with our theoretical analysis, let us establish some general assumptions.
-Based on our data generation model, we can only identify the left singular space and the singular values of $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$.
-Thus, without loss of generality, we can assume that the singular value decomposition of $\mathbf{V}_{M}^{\star}$ is $\mathbf{V}_{M}^{\star}=\mathbf{U}_{M}^{\star}\bm{\Lambda}_{M}^{\star}$ for $M\in\{1,2\}$, where $\mathbf{U}_{M}^{\star}\in\mathcal{O}_{d_{M},p}$ represents the left singular vectors, and $\bm{\Lambda}_{M}^{\star}\in\mathbb{R}^{p\times p}$ is a diagonal matrix containing the corresponding singular values.
-The embedding can only be identified up to a mean shift.
-For simplicity, we assume $\mathbf{V}_{M}^{\star\top}\bm{1}_{d_{M}}=\bm{0}$.
+In this section, we establish the theoretical properties of the proposed CLAIME algorithm and provide a comparative analysis with the two representative baselines discussed in Section 2.3: the simple concatenation approach (Concate) and conventional contrastive learning (CL).
+Our goal is to articulate the statistical principles that underlie CLAIME, thereby clarifying both why and how it achieves performance gains relative to these alternatives.
+3.1 Theoretical Properties of CLAIME
+First, we introduce some basic definitions.
+For $w\in{\cal W}^{(M)}$ and $t\in[T_{i}^{(M)}]$, define $X_{i,w}^{(M)}(t)=\mathbb{I}\{w^{(M)}_{i,t}=w\}$.
+According to model (2.1), conditioned on $\bm{c}_{i}$ and ${\bm{\epsilon}}_{i}^{(M)}=(\epsilon_{i,w}^{(M)})_{w\in{\cal W}^{(M)}}$, the variable $X_{i,w}^{(M)}(t)$ follows a Bernoulli distribution with mean
+$p_{i,w}^{(M)}=\mathbb{E}[X_{i,w}^{(M)}(t)\mid\bm{c}_{i},{\bm{\epsilon}}_{i}^{(M)}]=\frac{\exp(\langle\mathbf{v}_{w}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w}^{(M)})}{\sum_{w^{\prime}\in{\cal W}^{(M)}}\exp(\langle\mathbf{v}_{w^{\prime}}^{\star},\bm{c}_{i}\rangle+\epsilon_{i,w^{\prime}}^{(M)})}.$ (3.1)
+The probability $p_{i,w}^{(M)}$ depends on the discourse vector $\bm{c}_{i}$ and the noise vector ${\bm{\epsilon}}_{i}^{(M)}$, and is independent of $t$. Let $p_{w}^{(M)}:=\mathbb{E}[p_{i,w}^{(M)}]$.
+Next, define the co-occurrence of $w\in{\cal W}^{(M)}$ at time $t$ for patient $i$ and $w^{\prime}\in{\cal W}^{(M^{\prime})}$ at time $s$ for patient $j$ as $X_{i,j,w,w^{\prime}}^{(M,M^{\prime})}(t,s)=\mathbb{I}\{w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}$ for $t\neq s$, with mean $p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}$.
+When $i=j$, define $p_{w,w^{\prime}}^{(M,M^{\prime})}:=\mathbb{E}[p_{i,i,w,w^{\prime}}^{(M,M^{\prime})}]$.
+For $i\neq j$, we have $\mathbb{E}[p_{i,j,w,w^{\prime}}^{(M,M^{\prime})}]=p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}$.
+We then define the population PMI matrix $\mathbb{PMI}$ as
+$\mathbb{PMI}:=\begin{pmatrix}\mathbb{PMI}^{(1,1)}&\mathbb{PMI}^{(1,2)}\\ \mathbb{PMI}^{(2,1)}&\mathbb{PMI}^{(2,2)}\end{pmatrix},\quad\mathbb{PMI}^{(M,M^{\prime})}(w,w^{\prime})=\log\frac{p_{w,w^{\prime}}^{(M,M^{\prime})}}{p_{w}^{(M)}p_{w^{\prime}}^{(M^{\prime})}}$ (3.2)
+for $w\in{\cal W}^{(M)}$, $w^{\prime}\in{\cal W}^{(M^{\prime})}$, and $M,M^{\prime}\in\{1,2\}$.
+Before proceeding, we clarify our identifiability assumptions.
+From (2.1), only the left singular space and singular values of $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$ are identifiable.
+Hence, without loss of generality, we assume $\mathbf{V}_{M}^{\star}=\mathbf{U}_{M}^{\star}\bm{\Lambda}_{M}^{\star}$ for $M\in\{1,2\}$, where $\mathbf{U}_{M}^{\star}\in\mathcal{O}_{d_{M},p}$ contains the left singular vectors and $\bm{\Lambda}_{M}^{\star}\in\mathbb{R}^{p\times p}$ is diagonal.
+Model (2.1) is also invariant under the shift $\mathbf{v}_{w}^{\star}\leftarrow\mathbf{v}_{w}^{\star}-(1/d_{M})\sum_{w^{\prime}\in{\cal W}^{(M)}}\mathbf{v}_{w^{\prime}}^{\star}$ for all $w\in{\cal W}^{(M)}$, and ${\bm{\Sigma}}_{M}\leftarrow(\mathbf{I}-(1/d_{M})\bm{1}_{d_{M}}\bm{1}_{d_{M}}^{\scriptscriptstyle\sf T}){\bm{\Sigma}}_{M}$.
+Hence, we may also assume $\mathbf{V}_{M}^{\star{\scriptscriptstyle\sf T}}\bm{1}_{d_{M}}=\bm{0}$ and ${\bm{\Sigma}}_{M}^{\scriptscriptstyle\sf T}\bm{1}_{d_{M}}=\bm{0}$.
 Assumption 3.1.
 For $M\in\{1,2\}$, assume that $\|\mathbf{U}_{M}^{\star}\|_{2,\infty}\lesssim\sqrt{p/d_{M}}$, $\|\bm{\Lambda}_{M}^{\star}\|/s_{p}(\bm{\Lambda}_{M}^{\star})\lesssim 1$, and $\|\bm{\Lambda}_{M}^{\star}\|\ll 1$.
-Assumption 3.1 is widely known as the incoherence constant condition in the literature appearing in matrix completion (Candès and Recht,, 2009), PCA (Zhang et al.,, 2022), and the analysis of representation learning algorithm (Ji et al.,, 2021).
-We introduce the notation $\underline{d}=\min(d_{1},d_{2})$, and ${\bm{\Sigma}}=\mathrm{diag}({\bm{\Sigma}}_{1},{\bm{\Sigma}}_{2})$.
+Assumption 3.1 is widely known as the incoherence constant condition in the literature appearing in matrix completion (Candès and Recht, , 2009), PCA (Zhang et al., , 2022), and the analysis of representation learning algorithm (Ji et al., , 2023).
 Assumption 3.2.
-Assume that $\log d\ll d_{1}\wedge d_{2}$, $p\log^{12}d\ll d_{1}\wedge d_{2}$, $p\ll(\log\log d)^{2}$, $n\gg p^{2}d^{5}\log^{2}d$ and $p^{3}\log^{6}d\ll\underline{d}^{3/2}/d^{1/2}$.
-Assumption 3.2 imposes a condition on the dimension of the low-rank representation.
-This assumption is technically necessary in bounding the normalizing constant of the log-linear word production model (Lemma S5.2).
-Although, we can relax the assumption of $p\ll(\log\log d)^{2}$ into $p\ll\log^{2}d$, we stick to the current assumption for the sake of clarity.
+Assume that $d_{1}\wedge d_{2}\gg p\log^{2}(n+d)$ and $(d_{1}\wedge d_{2})^{2}\gg p^{4}(d_{1}\vee d_{2})\log(n+d)$, $n\gg p^{2}(d_{1}^{2}\vee d_{2}^{2})\log^{2}(n+d)$, and $n(S_{1}^{(1)}\wedge S_{1}^{(2)})\gg p^{2}(d_{1}^{6}\vee d_{2}^{6})\log^{2}(n+d)$.
+Assumption 3.2 imposes a condition on the regime of unique features $d_{1}$ and $d_{2}$, and the number of patients $n$.
+We will discuss the sufficiency of the condition in Remark 3.1.
+This assumption is technically necessary in bounding the normalizing constant of the model (Lemma S5.2).
+We introduce the notation ${\bm{\Sigma}}=\mathrm{diag}({\bm{\Sigma}}_{1},{\bm{\Sigma}}_{2})$.
 Assumption 3.3.
@@
 Assumption 3.3 implies that the signal-to-noise ratio is bounded above and below.
-Note that the signal, measured by the smallest positive singular value of $(1/p)\mathbf{V}_{M}^{\star}\mathbf{V}_{M^{\prime}}^{\star\top}$ for $M,M^{\prime}\in\{1,2\}$ is of order $1/p$ from Assumption 3.1.
+Note that the signal, measured by the smallest positive singular value of $(1/p)\mathbf{V}_{M}^{\star}\mathbf{V}_{M^{\prime}}^{\star{\scriptscriptstyle\sf T}}$ for $M,M^{\prime}\in\{1,2\}$ is of order $1/p$ from Assumption 3.1.
 Theorem 3.1.
 Under Assumptions 3.1, 3.2 and 3.3, we have
-$\Big{\|}\mathbb{PMI}^{(M,M)}-\quantity(\frac{1}{p}\mathbf{V}_{M}^{\star}\mathbf{V}_{M}^{\star\top}+{\bm{\Sigma}}_{M})\Big{\|}$ $\lesssim\frac{p^{3/2}}{d_{M}}\log^{6}d,\ \ M\in\{1,2\}$
-$\Big{\|}\mathbb{PMI}^{(1,2)}-\frac{1}{p}\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star\top}\Big{\|}$ $\lesssim\frac{p^{3/2}}{\underline{d}}\log^{6}d,$
-and hence
-$\Big{\|}\mathbb{PMI}-\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}})\Big{\|}$ $\lesssim\frac{p^{2}}{\underline{d}}\log^{6}d.$
+$\Big{\|}\mathbb{PMI}^{(M,M)}-\quantity(\frac{1}{p}\mathbf{V}_{M}^{\star}\mathbf{V}_{M}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}}_{M})\Big{\|}_{\mathrm{F}}$ $\lesssim\frac{p^{2}}{d_{M}}\log(n+d),\ \ M\in\{1,2\}$
+$\Big{\|}\mathbb{PMI}^{(1,2)}-\frac{1}{p}\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star{\scriptscriptstyle\sf T}}\Big{\|}_{\mathrm{F}}$ $\lesssim\frac{p^{2}(d_{1}\vee d_{2})}{(d_{1}\wedge d_{2})^{2}}\log(n+d),\quad\text{ and hence }$
+$\Big{\|}\mathbb{PMI}-\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}})\Big{\|}_{\mathrm{F}}$ $\lesssim\frac{p^{2}(d_{1}\vee d_{2})}{(d_{1}\wedge d_{2})^{2}}\log(n+d).$
 The proof of Theorem 3.1 is presented in Supplementary S4.1.
-This theorem demonstrates that the PMI matrix for each modality can be closely approximated by the embedding space plus the noise covariance matrix ${\bm{\Sigma}}_{M}$.
+It shows that the PMI matrix for each modality can be closely approximated by the product of embeddings plus the noise covariance matrix ${\bm{\Sigma}}_{M}$.
 When the noise level ${\bm{\Sigma}}_{M}$ is substantial, the PMI value $\mathbb{PMI}^{(M,M)}(w,w^{\prime})$ for a single modality may significantly deviate from $\langle\mathbf{v}_{w}^{\star},\mathbf{v}_{w^{\prime}}^{\star}\rangle/p$.
-On the other hand, the cross-modal PMI matrix $\mathbb{PMI}^{(1,2)}$ can be directly approximated by $\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star\top}/p$.
-The theorem below affirms that $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an effective estimator for $\mathbb{PMI}^{(1,2)}$.
+On the other hand, the cross-modal PMI matrix $\mathbb{PMI}^{(1,2)}$ can be directly approximated by $\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star{\scriptscriptstyle\sf T}}/p$.
+The next theorem affirms that $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an effective estimator for $\mathbb{PMI}^{(1,2)}$.
+Assumption 3.4.
+For $M=1,2$, assume that $S_{2}^{(M)2}\gg S_{1}^{(M)}$, $S_{1}^{(M)}\gtrsim S_{2}^{(M)}\gtrsim S_{4}^{(M)}$ and $S_{2}^{(1,2)}\gtrsim S_{4}^{(1,2)}$.
+Also assume the non-negative correlation between $(T_{i}^{(1)})_{i\in[n]}$ and $(T_{i}^{(2)})_{i\in[n]}$, i.e., $S_{2}^{(1,2)2}\geq S_{1}^{(1)}S_{1}^{(2)}$.
+Assumption 3.4 requires that the empirical variance of $T_{i}^{(M)}$ dominates the mean of $T_{i}^{(M)}$, and their moments are comparable to each other.
+For the same length setting $T_{i}^{(M)}\equiv T^{(M)}$ for all $i$, the assumption is satisfied if $T^{(M)}\to\infty$.
+Assumption 3.4 is satisfied for exponential distribution and gamma distribution with constant order shape parameter.
 Theorem 3.2.
-Suppose that $n\gg d^{2}\log^{2}d$. Under Assumptions 3.1-3.3 and 3.5, we have
-$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}-\mathbb{PMI}^{(1,2)}}_{\max}$ $\lesssim\frac{d_{1}d_{2}}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}^{3/2}}\log^{6}d\ \ \text{ and }$
-$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}-\mathbb{PMI}^{(1,2)}}$ $\lesssim\frac{\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}}\log^{6}d$
-with probability $1-\exp(-\Omega(\log^{2}d))$.
+Under Assumptions 3.1-3.3 and 3.4, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
+$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}-\mathbb{PMI}^{(1,2)}}_{\mathrm{F}}$ $\lesssim\left\{\frac{d_{1}^{3/2}d_{2}^{3/2}}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{d_{1}^{1/2}d_{2}^{1/2}}{\sqrt{n}}+\frac{d_{1}^{1/2}d_{2}^{1/2}p^{2}}{(d_{1}\wedge d_{2})^{2}}\right\}\log(n+d).$
 The proof of Theorem 3.2 is given in Supplementary S4.3.
-By Assumption 3.2, it implies that $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an asymptotically unbiased estimate of $\mathbb{PMI}^{(1,2)}$.
-Combining Theorems 3.1 and 3.2, we get the rate of CLAIME.
-We also provide a rate of learning dual representations jointly by CLAIME.
-Namely, we consider concatenating the left singular vectors of two matrices $\mathbf{\widehat{V}}_{1}$ and $\mathbf{\widehat{V}}_{2}$.
-Since $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\top})$ are only identifiable up to rotation, the subspace spanned by the column vectors of $\mathbf{P}_{p}([\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{2})]$ depends on the choice of rotations.
-Hence we introduce a matrix $\mathbf{H}$, and define the estimated joint representation by $\mathbf{\widehat{U}}_{\mathbf{H}}:=\mathbf{P}_{p}([\mathbf{H}\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{P}_{p}^{\top}(\mathbf{\widehat{V}}_{2}^{\top})])$, where $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\top})$ and $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\top})$ are arbitrary top-$p$ left singular vectors of $\mathbf{\widehat{V}}_{1}$ and $\mathbf{\widehat{V}}_{2}$, respectively.
-We then measure the best performance of $\mathbf{\widehat{U}}_{\mathbf{H}}$ with respect to $\mathbf{H}$ by $\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$, where $\mathbf{U}^{\star}\in\mathcal{O}_{d,p}$ is the concatenated embedding vectors defined as $\mathbf{U}^{\star}:=\mathbf{P}_{p}([\mathbf{U}_{1}^{\star\top},\mathbf{U}_{2}^{\star\top}])$.
+Under Assumption 3.2, $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ is an asymptotically unbiased estimate of $\mathbb{PMI}^{(1,2)}$.
+Combining Theorems 3.1 and 3.2 yields the convergence rate of CLAIME.
+We also analyze the rate of learning joint representations via CLAIME.
+Specifically, we concatenate the left singular vectors of $\mathbf{\widehat{V}}_{1}$ and $\mathbf{\widehat{V}}_{2}$.
+Since $\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\scriptscriptstyle\sf T})$ are only identifiable up to rotation, the subspace spanned by $\mathbf{P}_{p}([\mathbf{P}_{p}^{\scriptscriptstyle\sf T}(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}^{\scriptscriptstyle\sf T}(\mathbf{\widehat{V}}_{2})]$ depends on the choice of rotations.
+To address this, we introduce a matrix $\mathbf{H}$ and define the estimated joint representation by $\mathbf{\widehat{U}}_{\mathbf{H}}:=\mathbf{P}_{p}([\mathbf{H}\mathbf{P}_{p}^{\scriptscriptstyle\sf T}(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}^{\scriptscriptstyle\sf T}(\mathbf{\widehat{V}}_{2}^{\scriptscriptstyle\sf T})])$.
+We then evaluate performance via $\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$, where $\mathbf{U}^{\star}\in\mathcal{O}_{d,p}$ is the concatenated embedding vectors defined as $\mathbf{U}^{\star}:=(1/\sqrt{2})[\mathbf{U}_{1}^{\star{\scriptscriptstyle\sf T}},\mathbf{U}_{2}^{\star{\scriptscriptstyle\sf T}}]$.
+The proof of Theorem 3.3 is presented in Supplementary S4.3.
+Theorem $3.$
+Under Assumptions 3.1-3.4, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
+$\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T}),\mathbf{U}_{1}^{\star})\|_{\mathrm{F}}\vee\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\scriptscriptstyle\sf T}),\mathbf{U}_{2}^{\star})\|_{\mathrm{F}}\lesssim\frac{p^{3}(d_{1}\vee d_{2})\log(n+d)}{(d_{1}\wedge d_{2})^{2}}$
+$\quad\quad+\frac{pd_{1}^{3/2}d_{2}^{3/2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{pd_{1}^{1/2}d_{2}^{1/2}\log(n+d)}{\sqrt{n}}\quad\text{ and }$ (3.3)
+$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\left\{\frac{p^{3}(d_{1}\vee d_{2})}{(d_{1}\wedge d_{2})^{2}}+\frac{pd_{1}^{3/2}d_{2}^{3/2}}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{pd_{1}^{1/2}d_{2}^{1/2}}{\sqrt{n}}\right\}\log(n+d).$
+Remark 3.1.
+The bound in (3.3) consists of a bias term from the nonlinearity of the data-generating process (first term), and two variance terms (last two terms).
+The first variance term arises from estimating per-record co-occurrence frequencies.
+Each patient contributes $(T_{i}^{(1)},T_{i}^{(2)})$-length multimodal data, so the variance of empirical co-occurrence frequencies is $O(1/(n(S_{1}^{(1)}\wedge S_{1}^{(2)})))$.
+Normalizing to obtain PMI entries scales this variance by $d_{1}d_{2}$, and converting to Frobenius norm adds an extra factor of $d_{1}^{1/2}d_{2}^{1/2}$ via the inequality $\|A\|_{\mathrm{F}}\leq d_{1}^{1/2}d_{2}^{1/2}\|A\|_{\max}$.
+Applying the Davis–Kahan theorem then gives a contribution of order $O(pd_{1}^{3/2}d_{2}^{3/2}/(n(S_{1}^{(1)}\wedge S_{1}^{(2)})))$.
+The second variance term reflects the averaging across $n$ patients, yielding a $1/\sqrt{n}$ rate, which again is scaled by $d_{1}^{1/2}d_{2}^{1/2}$ when converted to Frobenius norm.
+As a result, to make the overall bound vanish and ensure consistency, Assumption 3.2 is required.
+3.2 Comparison with the baseline methods
+We now compare CLAIMEwith the two baseline approaches introduced in Section 2.3.
+As discussed there, both baselines can be interpreted as applying SVD to different empirical PMI matrices: the first is SVD on the concatenated PMI matrix $\widehat{\mathbb{PMI}}$, while the second is based on the estimator $\widehat{\mathbb{PMI}}_{\textnormal{CL}}$ derived from the CL loss (2.5).
+We show in Theorem 3.4 (proof in Supplementary S4.4) that both $\widehat{\mathbb{PMI}}$ and $\widehat{\mathbb{PMI}}_{\textnormal{CL}}$ are biased estimators of the population matrix $\mathbb{PMI}$.
+Building on this result, Theorem 3.5 establishes the convergence rates of the corresponding estimators $\widehat{\mathbf{V}}_{\textnormal{CL}}$ and $\mathbf{\widehat{V}}_{\textnormal{Con}}$, with detailed proofs provided in Supplementary S4.4.
 Theorem 3.4.
-Suppose that $n\gg d^{2}\log^{2}d$. Under Assumptions 3.1-3.3 and 3.5,
-$\norm{\mathbf{\widehat{V}}_{1}\mathbf{\widehat{V}}_{2}^{\top}-\frac{1}{\lambda p}\mathbf{V}_{1}^{\star}\mathbf{V}_{2}^{\star\top}}\lesssim\frac{1}{\lambda}\quantity(\frac{\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}}\log^{6}d),$
-$\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{1}^{\top}),\mathbf{U}_{1}^{\star})\|_{\mathrm{F}}\vee\|\sin\Theta(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{2}^{\top}),\mathbf{U}_{2}^{\star})\|_{\mathrm{F}}\lesssim\frac{p^{3/2}\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d$
-and
-$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\frac{p^{3/2}\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d$
-with probability $1-\exp(-\Omega(\log^{2}d))$.
-The proof of Theorem 3.3 is presented in Supplementary S4.3.
-On the contrary, we can show that the empirical concatenated PMI matrix $\widehat{\mathbb{PMI}}$ and $\widehat{\mathbb{PMI}}_{\textnormal{CL}}$ derived from the CL loss (2.6) are biased estimators for $\mathbb{PMI}$.
+Under Assumptions 3.1-3.3 and 3.4, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
+$\norm{\widehat{\mathbb{PMI}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{Con}})}_{\mathrm{F}}$ $\lesssim\frac{(d_{1}\vee d_{2})d_{1}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(1)1/2}}+\frac{(d_{1}\vee d_{2})d_{2}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(2)1/2}}$
+$\quad+\frac{(d_{1}\vee d_{2})d_{1}d_{2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{(d_{1}\vee d_{2})\log(n+d)}{\sqrt{n}},\quad\text{ and}$
+$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CL}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{CL}})}_{\mathrm{F}}\lesssim\frac{(d_{1}\vee d_{2})d_{1}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(1)1/2}}+\frac{(d_{1}\vee d_{2})d_{2}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(2)1/2}}$
+$\quad\quad+\frac{(d_{1}\vee d_{2})d_{1}d_{2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{\log(n+d)(d_{1}\vee d_{2})}{\sqrt{n}}+\frac{p^{2}(d_{1}\vee d_{2})\log(n+d)}{(d_{1}\wedge d_{2})^{2}}.$
+Here, $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ are bias matrices defined in Supplementary S4.2 and S4.4.
+Remark 3.2.
+As shown in Theorems S4.2 (Concate) and S4.5 (CL), the bias matrix $\mathbf{B}_{\textnormal{CL}}$ includes a correctable term based on observed $T_{i}^{(M)}$ and an uncorrectable term involving the population-level PMI.
+In contrast, $\mathbf{B}_{\textnormal{Con}}$ depends only on $T_{i}^{(M)}$ and is fully correctable.
+To evaluate the effect of this correction, we conducted additional experiments by removing the bias term associated with $T_{i}^{(M)}$.
+The results, presented in Supplementary S7.3.3, show that the debiased estimators offer little to no improvement over the original ones, suggesting that the impact of the removable bias component is limited in practice.
 Theorem 3.5.
-Suppose that $n\gg d^{2}\log^{2}d$. Under Assumptions 3.1-3.3 and 3.5, we have
-$\norm{\widehat{\mathbb{PMI}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{Con}})}_{\max}\lesssim\frac{d^{2}}{\sqrt{n}}\log d,\quad\norm{\widehat{\mathbb{PMI}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{Con}})}\lesssim\frac{d^{5/2}}{\sqrt{n}}\log d,$
-and
-$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CL}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{CL}})}_{\max}$ $\lesssim\frac{d^{2}}{\sqrt{n}}\log d+\frac{p^{3/2}}{\underline{d}^{3/2}}\log^{6}d,$
-$\norm{\widehat{\mathbb{PMI}}_{\textnormal{CL}}-(\mathbb{PMI}+\mathbf{B}_{\textnormal{CL}})}$ $\lesssim\frac{d^{5/2}}{\sqrt{n}}\log d+\quantity(\frac{d}{\underline{d}})^{1/2}\frac{p^{3/2}}{\underline{d}}\log^{6}d.$
-Here the matrices $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ reflect the bias terms defined in Supplementary S4.2 and S4.4.
-Remark 3.1.
-The bias matrices $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ depend only on $T_{i}^{(M)}$’s, which can be effectively removed since $T_{i}^{(M)}$’s are directly observed.
-The proof of Theorem 3.4 is presented in Supplementary S4.4.
-Based on this theorem, we are ready to show the rate of $\widehat{\mathbf{V}}_{\textnormal{CL}}$ and $\mathbf{\widehat{V}}_{\textnormal{Con}}$.
-Before that, we need the following assumption.
-Assumption 3.4.
-Assume that
-$\frac{s_{p}({\bm{\Sigma}})-s_{p+1}({\bm{\Sigma}})}{s_{p}({\bm{\Sigma}})}\gtrsim 1.$
-Assumption 3.4 is known as the eigengap condition for the top-$p$ singular vectors of ${\bm{\Sigma}}$.
-A similar assumption has been used in the analysis of contrastive learning (Ji et al.,, 2021).
-This enables us to identify the top-$p$ singular vectors of $(1/p)\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}}$, resulting in the following theorem.
-Theorem 3.5.
-Suppose that $n\gg d^{2}\log^{2}d$ and $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1/p$. Under Assumptions 3.1-3.5,
-$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{\textnormal{Con}}^{\top}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{Con}}))}_{\mathrm{F}}$ $\lesssim\frac{p^{3/2}d^{5/2}}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d,$
-$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\widehat{\mathbf{V}}_{\textnormal{CL}}^{\top}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star\top}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{CL}}))}_{\mathrm{F}}$ $\lesssim\frac{p^{3}d^{5/2}}{\sqrt{n}}\log d+\quantity(\frac{d}{\underline{d}})^{1/2}\frac{p^{9/2}}{\underline{d}}\log^{6}d$
-holds with probability $1-\exp(-\Omega(\log^{2}d))$.
-The proof Theorem 3.5 is given in Supplementary S4.4.
-It shows that both Concate and CL will yield biased estimators for the embeddings $\mathbf{V}^{\star}$ due to the inherent bias terms and covariance matrix of the noise.
-While the bias terms $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ can be removed, the influence of ${\bm{\Sigma}}$ remains irreducible due to its unobserved nature.
+Suppose that $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1/p$. Under Assumptions 3.1-3.4, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
+$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\mathbf{\widehat{V}}_{\textnormal{Con}}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{Con}}))}_{\mathrm{F}}$
+$\quad\lesssim\frac{p(d_{1}\vee d_{2})d_{1}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(1)1/2}}+\frac{p(d_{1}\vee d_{2})d_{2}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(2)1/2}}$
+$\quad\quad+\frac{p(d_{1}\vee d_{2})d_{1}d_{2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{p\log(n+d)(d_{1}\vee d_{2})}{\sqrt{n}}+\frac{p^{3}(d_{1}\vee d_{2})\log(n+d)}{(d_{1}\wedge d_{2})^{2}},$
+$\norm{\sin\Theta\quantity(\mathbf{P}_{p}(\widehat{\mathbf{V}}_{\textnormal{CL}}^{\scriptscriptstyle\sf T}),\mathbf{P}_{p}\quantity(\frac{1}{p}\mathbf{V}^{\star}\mathbf{V}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}}+\mathbf{B}_{\textnormal{CL}}))}_{\mathrm{F}}$
+$\quad\lesssim\frac{p(d_{1}\vee d_{2})d_{1}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(1)1/2}}+\frac{p(d_{1}\vee d_{2})d_{2}^{2}\log(n+d)}{\sqrt{n}S_{1}^{(2)1/2}}$
+$\quad\quad+\frac{p(d_{1}\vee d_{2})d_{1}d_{2}\log(n+d)}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{p\log(n+d)(d_{1}\vee d_{2})}{\sqrt{n}}+\frac{p^{3}(d_{1}\vee d_{2})\log(n+d)}{(d_{1}\wedge d_{2})^{2}}.$
+Theorem 3.5 shows that both Concate and CL will yield biased estimators for the embeddings $\mathbf{V}^{\star}$ due to the inherent bias terms and covariance matrix of the noise.
+While the bias terms $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$ can be partially removed, the influence of $\bm{\Sigma}$} remains irreducible due to its unobserved nature.
 Comparing Theorems 3.3 and 3.5, we find that CLAIME achieves a sharper rate than Concate and CL even when $\mathbf{B}_{\textnormal{CL}}$, $\mathbf{B}_{\textnormal{Con}}$ and $\bm{\Sigma}$} go to zero.
 As a result, CLAIME is a better estimator than Concate and CL.
-To quantify the effect of $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$, we define the following empirical moments of $T_{i}^{(M)}$: $S_{q}^{(M)}:=((1/n)\sum_{i\in[n]}T_{i}^{(M)q})^{1/q}$.
-We also define the empirical cross-moments as $S_{q}^{(1,2)}:=((1/n)\sum_{i\in[n]}(T_{i}^{(1)})^{q/2}(T_{i}^{(2)})^{q/2})^{1/q}$ for $q=2,4$.
-Note that when the length of patient data is the same, i.e., $T_{i}^{(M)}=T^{(M)}$ for all $i\in[n]$ for $M=1,2$, $S_{q}^{(M)}=T^{(M)}$ for all $q\geq 1$.
-Assumption 3.5.
-For $M=1,2$, assume that $1\gtrsim S_{1}^{(1)}S_{1}^{(2)}/S_{2}^{(1,2)2}\gg 1/n$, $S_{1}^{(M)2}/S_{2}^{(M)2}\gg 1/n$ and $S_{2}^{(M)2}\gg S_{1}^{(M)}$.
-Also assume that $S_{1}^{(M)}\gtrsim S_{2}^{(M)}\gtrsim S_{4}^{(M)}$ and $S_{2}^{(1,2)}\gtrsim S_{4}^{(1,2)}$.
-Assumption 3.5 requires that the empirical variance of $T_{i}^{(M)}$ dominates the mean of $T_{i}^{(M)}$, and their moments are comparable to each other.
-For the same length setting $T_{i}^{(M)}\equiv T^{(M)}$ for all $i\in[n]$, the assumption is satisfied if $T^{(M)}\to\infty$ and $n\to\infty$.
-For the bias term $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$, we have the following lemma.
+For the bias terms $\mathbf{B}_{\textnormal{Con}}$ and $\mathbf{B}_{\textnormal{CL}}$, we have the following lemma.
 The proof of Lemma 3.1 is given in Supplementary S4.7.
 Lemma 3.1.
-Suppose that Assumption 3.5 holds. If $T_{i}^{(1)}=T_{i}^{(2)}$, then,
-$\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim d\frac{S_{1}^{(1)}}{S_{2}^{(1)2}}.$
-Furthermore, if $T_{i}^{(1)}=T_{i}^{(2)}\equiv T$ for all $i\in[n]$ for some $T>0$, then
-$\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim\frac{d}{T}.$
-We then characterize the effect of ${\bm{\Sigma}}$.
-We will show that when ${\bm{\Sigma}}$ satisfies suitable conditions, the two existing methods Concate and CL will not be an ideal estimator of $\mathbf{V}^{\star}$.
+Suppose that Assumption 3.4 holds.
+If $T_{i}^{(1)}=T_{i}^{(2)}$, then, $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim(d_{1}\vee d_{2})S_{1}^{(1)}/S_{2}^{(1)2}$.
+Furthermore, if $T_{i}^{(1)}=T_{i}^{(2)}\equiv T$ for all $i\in[n]$ for some $T>0$, then $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\lesssim(d_{1}\vee d_{2})/T$.
+We then characterize the effect of $\bm{\Sigma}$} and show that when $\bm{\Sigma}$} satisfies suitable conditions, the two existing methods Concate and CL will not provide ideal estimators of $\mathbf{V}^{\star}$.
+Assumption 3.5.
+Assume that $(s_{p}({\bm{\Sigma}})-s_{p+1}({\bm{\Sigma}}))/s_{p}({\bm{\Sigma}})\gtrsim 1$.
+Assumption 3.5 is known as the eigengap condition for the top-$p$ singular vectors of $\bm{\Sigma}$}.
+A similar assumption has been used in the analysis of contrastive learning (Ji et al., , 2023).
+This enables us to identify the top-$p$ singular vectors of $(1/p)\mathbf{V}^{\star}\mathbf{V}^{\star{\scriptscriptstyle\sf T}}+{\bm{\Sigma}}$.
 Assumption 3.6.
@@
 Theorem 3.6.
-Suppose that Assumptions 3.1-3.6 hold. If $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1$, then,
-$\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{Con}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\gtrsim\sqrt{p}\ \ \text{ and }\ \ \|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{CL}},\mathbf{U}^{\star})\|_{\mathrm{F}}\gtrsim\sqrt{p}$
-with probability $1-\exp(-\Omega(\log^{2}d))$.
+Suppose $\|\mathbf{B}_{\textnormal{Con}}\|\vee\|\mathbf{B}_{\textnormal{CL}}\|\ll 1/p$.
+Under Assumptions 3.1-3.6, with probability $1-\exp(-\Omega(\log^{2}(n+d)))$, $\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{Con}},\mathbf{U}^{\star})\|_{\mathrm{F}}\gtrsim\sqrt{p}$ and $\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{CL}},\mathbf{U}^{\star})\|_{\mathrm{F}}\gtrsim\sqrt{p}$.
 The formal statement is available in Theorem S4.8.
@@
 Corollary 3.1.
-Choose $\mathbf{P}^{(1)}\in\mathcal{O}_{d_{1},p}$ satisfying $[\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\in\mathcal{O}_{d_{1},2p}$ and $\|[\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\|_{2,\infty}^{2}\lesssim p/d_{1}$.
-Similarly choose $\mathbf{P}^{(2)}\in\mathcal{O}_{d_{2},p}$ such that $[\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\in\mathcal{O}_{d_{2},2p}$ and $\|[\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\|_{2,\infty}^{2}\lesssim p/d_{2}$.
-Let ${\bm{\Sigma}}_{1}:=\mathbf{P}^{(1)}\mathbf{P}^{(1)\top}$, ${\bm{\Sigma}}_{2}:=(1/2)\mathbf{P}^{(2)}\mathbf{P}^{(2)\top}$, $\mathbf{V}^{\star}_{1}=\mathbf{U}^{\star}_{1}$ and $\mathbf{V}^{\star}_{2}=\mathbf{U}^{\star}_{2}$.
-For these noise covariance matrices, Assumptions 3.1, 3.3 and 3.4 hold.
-In addition, Assumption 3.6 holds since $\mathbf{P}^{(1)\top}\mathbf{V}_{1}^{\star}=\mathbf{P}^{(2)\top}\mathbf{V}_{2}^{\star}=\mathbf{O}_{p\times p}$.
-Under additional assumptions on $(T_{i}^{(M)})_{i\in[n]}$, $d_{1}$ and $d_{2}$ (Assumptions 3.2 and 3.5), we obtain
+Choose $\mathbf{P}^{(1)}\in\mathcal{O}_{d_{1},p}$ satisfying $\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\in\mathcal{O}_{d_{1},2p}$ and $\|[\mathbf{U}_{1}^{\star};\mathbf{P}^{(1)}]\|_{2,\infty}^{2}\lesssim p/d_{1}$.
+Similarly choose $\mathbf{P}^{(2)}\in\mathcal{O}_{d_{2},p}$ such that $\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\in\mathcal{O}_{d_{2},2p}$ and $\|[\mathbf{U}_{2}^{\star};\mathbf{P}^{(2)}]\|_{2,\infty}^{2}\lesssim p/d_{2}$.
+Let ${\bm{\Sigma}}_{1}:=\mathbf{P}^{(1)}\mathbf{P}^{(1){\scriptscriptstyle\sf T}}$, ${\bm{\Sigma}}_{2}:=(1/2)\mathbf{P}^{(2)}\mathbf{P}^{(2){\scriptscriptstyle\sf T}}$, $\mathbf{V}^{\star}_{1}=\mathbf{U}^{\star}_{1}$ and $\mathbf{V}^{\star}_{2}=\mathbf{U}^{\star}_{2}$.
+For these noise covariance matrices, Assumptions 3.1, 3.3 and 3.5 hold.
+In addition, Assumption 3.6 holds since $\mathbf{P}^{(1){\scriptscriptstyle\sf T}}\mathbf{V}_{1}^{\star}=\mathbf{P}^{(2){\scriptscriptstyle\sf T}}\mathbf{V}_{2}^{\star}=\mathbf{O}_{p\times p}$.
+Under additional assumptions on $(T_{i}^{(M)})_{i\in[n]}$, $d_{1}$ and $d_{2}$ (Assumptions 3.2 and 3.4), with probability $1-\exp(-\Omega(\log^{2}(n+d)))$,
 $\|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{Con}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\gtrsim\sqrt{p},\ \ \|\sin\Theta(\mathbf{\widehat{U}}_{\textnormal{CL}},\mathbf{U}^{\star})\|_{\mathrm{F}}\gtrsim\sqrt{p}\ \ \text{ and }$
-$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\frac{p^{3/2}\underline{d}^{3/2}d}{\sqrt{n}}\log d+\frac{p^{3}}{\underline{d}}\log^{6}d.$
+$\min_{\mathbf{H}\in\mathcal{O}_{p,p}}\|\sin\Theta(\mathbf{\widehat{U}}_{\mathbf{H}},\mathbf{U}^{\star})\|_{\mathrm{F}}$ $\lesssim\left\{\frac{p^{3}(d_{1}\vee d_{2})}{(d_{1}\wedge d_{2})^{2}}+\frac{pd_{1}^{3/2}d_{2}^{3/2}}{\sqrt{n}(S_{1}^{(1)1/2}\wedge S_{1}^{(2)1/2})}+\frac{pd_{1}^{1/2}d_{2}^{1/2}}{\sqrt{n}}\right\}\log(n+d).$
 Corollary 3.1 directly follows from Theorems 3.3 and 3.6.
 Note that the factor $1/2$ in the definition of ${\bm{\Sigma}}_{2}$ guarantees the eigengap condition at the $p$-th largest singular value of the joint noise covariance matrix $\bm{\Sigma}$}.
-The noise covariance in Corollary 3.1 implies that the subspace spanned by the noise is orthogonal to the subspace spanned by the signal, so that the noisy fluctuations in feature frequency distribution occurs simultaneously to many words, in a separable way from the signal.
-Under this condition, Concate and CL learn the eigenspace corresponding to noise covariance matrix, since those methods are vulnerable to large noise strength.
-On the other hand, CLAIME can effectively escape from the noise, and learns core representations when $n$ and $\bar{d}$ are sufficiently large.
+The noise covariance in Corollary 3.1 implies that the subspace spanned by the noise is orthogonal to that of the signal, so that noisy fluctuations in the feature frequency distribution occur simultaneously across many features, in a way separable from the signal.
+Under this condition, Concate and CL tend to learn the eigenspace corresponding to noise covariance matrix, as they are vulnerable to strong noise.
+In contrast, CLAIME effectively escapes from the noise and recovers the core representations when $n$ and $(d_{1}\vee d_{2})/(d_{1}\wedge d_{2})^{2}$ are sufficiently large.
 4 Simulation Studies
-In this section, we assess CLAIME’s performance and compare it to CL and Concate via simulation studies.
-When patient-level data is not available, these three methods can be deployed by performing SVD on different aggregated PMI matrices introduced in Section 2.
-Since patient-level data is also available in the synthesized datasets, we can additionally apply the gradient descent-based multimodal contrastive learning method (CLAIME-GD) with target function at (2.3), and the gradient descent-based contrastive learning method (CL-GD) with target function at (2.6), respectively.
+In this section, we evaluate the performance of CLAIME and compare it to CL, Concate, and Kernel PCA (KPCA) through simulation studies.
+KPCA is a nonlinear PCA method that operates in a high-dimensional feature space induced by a kernel function.
+Following Gupta et al., (2019), we apply KPCA with Gaussian RBF kernel directly to the co-occurrence matrix $\mathbf{D}$ given that the co-occurrence matrix can be viewed as a generalized $n$-gram similarity matrix.
+When patient-level data is unavailable, the first three methods can still be applied by performing SVD on various aggregated PMI matrices, as described in Section 2.
+Since patient-level data is also available in the synthesized datasets, we can additionally apply the gradient descent-based variants: CLAIME-GD and CL-GD, which optimize the objective functions defined in (2.3) and (2.5), respectively.
 The suffix “-GD” is added to distinguish them from their SVD-based counterparts.
@@
 Both gradient-based algorithms converge when their loss functions change by less than $10^{-6}$.
-We generate patient embedding vector $\mathbf{c}_{i}\in\mathbb{R}^{p}$ for each patient $i$ from a multivariate normal distribution $N(\mathbf{0},\mathbf{I}_{p})$ independently.
-We construct the code and CUI embedding matrices $\mathbf{V}_{1}^{\star}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}^{\star}\in\mathbb{R}^{d_{2}\times p}$, respectively, by generating each entry from a standard normal distribution.
-The row means of $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$ are standardized to zero, and their maximum singular values are standardized to $1$.
-Then, we follow the data generation process in (2.1) to generate $T_{i}^{(1)}$ number of codes and $T_{i}^{(2)}$ number of CUIs for the $i$-th patient, where $T_{i}^{(1)},T_{i}^{(2)}$ follow a Poisson distribution with the mean of $50$ and only non-zero values are retained.
-Two different cases of error covariance structures are considered.
-For Case 1, we set $\mathbf{\Sigma}^{(M)}=\text{diag}(\mathbf{I}_{p/2},\mathbf{0})/c\in\mathbb{R}^{d_{M}\times d_{M}}$ for $M=1$ or $2$, where $c>0$ is a constant controlling the signal-to-noise ratio (SNR) of the data generating process.
-For Case 2, we generated $\sigma_{w}\overset{i.i.d.}{\sim}\text{Unif}(0,1)$ and set ${\bm{\Sigma}}_{w,w^{\prime}}^{(M)}=\rho^{|w-w^{\prime}|}\sigma_{w}\sigma_{w^{\prime}}/2$ for $w,w^{\prime}\in[d_{M}]$ and $M\in\{1,2\}$, where $\rho\in(0,1)$ controls both the structure and the SNR of the data generating process.
-Specifically, a larger $\rho$ leads to a greater signal-to-noise ratio in the synthesized data.
-For both cases, three experiments are separately conducted to compare the performance of CLAIME, CLAIME-GD, CL, CL-GD, and Concate under different sample sizes $n$, feature sizes $d$, and SNR.
-Their performance is evaluated using the metric $\text{Err}(\widehat{\mathbf{V}},\mathbf{V}^{\star})=\max_{M=1,2}\|\widehat{\mathbf{V}}_{M}\widehat{\mathbf{V}}_{M}^{\top}-\mathbf{V}_{M}^{\star}\mathbf{V}_{M}^{\star\top}\|_{\mathrm{F}}$, which measures the distance between the estimated and true embedding subspaces.
+We generate a patient embedding vector $\mathbf{c}_{i}\in\mathbb{R}^{p}$ for each patient $i$, independently drawn from a multivariate normal distribution $N(\mathbf{0},\mathbf{I}_{p})$.
+To construct the code and CUI embedding matrices, we first generate $\mathbf{V}^{\star}=\mathbf{R}\mathbf{M}$, where $\mathbf{M}\in\mathbb{R}^{d\times p}$ has i.i.d.
+standard normal entries, and $\mathbf{R}\in\mathcal{O}_{d,d}$ is an orthonormal matrix obtained via QR decomposition of a random matrix.
+The rows of $\mathbf{V}^{\star}$ are mean-centered (i.e., each row has zero mean) and rescaled so that its largest singular value is $1$.
+We then partition $\mathbf{V}^{\star}$ into $\mathbf{V}_{1}^{\star}$ (first $d_{1}$ rows) and $\mathbf{V}_{2}^{\star}$ (remaining $d_{2}$ rows).
+Performing SVD on $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$ separately yields the column basis matrices $\mathbf{U}_{1}^{\star}$ and $\mathbf{U}_{2}^{\star}$.
+Next, we generate the data according to model (2.1), where the patient $i$ is assigned $T_{i}^{(1)}$ codes and $T_{i}^{(2)}$ CUIs, with $T_{i}^{(1)},T_{i}^{(2)}\sim\text{Poisson}(50)$ truncated to exclude zeros.
+We consider two different cases for the error covariance structure.
+For Case 1, we set $\mathbf{\Sigma}^{(M)}=\text{diag}(\mathbf{I}_{p/2},\mathbf{0})/c\in\mathbb{R}^{d_{M}\times d_{M}}$ for $M\in\{1,2\}$, where $c>0$ controls the signal-to-noise ratio (SNR) of the data generating process.
+For Case 2, we generated $\sigma_{w}\overset{i.i.d.}{\sim}\text{Unif}(0,1)$ and set ${\bm{\Sigma}}_{w,w^{\prime}}^{(M)}=\rho^{|w-w^{\prime}|}\sigma_{w}\sigma_{w^{\prime}}/2$ for $w,w^{\prime}\in[d_{M}]$ and $M\in\{1,2\}$, where $\rho\in(0,1)$ controls both the structure and the SNR.
+Specifically, larger $\rho$ values yield higher SNR in the synthesized data.
+For both cases, we set $d_{1}=d_{2}=d/2$ and conduct three experiments to compare the performance of CLAIME, CLAIME-GD, CL, CL-GD, Concate and KPCA under varying sample sizes $n$, feature sizes $d$, and SNR.
+Performance is evaluated using the metric $\text{Err}(\widehat{\mathbf{U}},\mathbf{U}^{\star})=\max_{M=1,2}\|\widehat{\mathbf{U}}_{M}\widehat{\mathbf{U}}_{M}^{{\scriptscriptstyle\sf T}}-\mathbf{U}_{M}^{\star}\mathbf{U}_{M}^{\star{\scriptscriptstyle\sf T}}\|_{\mathrm{F}}$, which measures the subspace distance between estimated and true embeddings.
 We first fix the SNR by setting $c=0.2$ for Case 1 and $\rho=0.8$ for Case 2.
-In the first experiment, we vary the sample sizes $n$ in $\{2\times 10^{4},4\times 10^{4},6\times 10^{4},8\times 10^{4},10^{5}\}$, while fixing $d=2d_{1}=2d_{2}=1000$ and $p=4$.
-In the second experiment, we vary $d$ in $\{200,400,600,800,1000\}$ and set $d_{1}=d_{2}=d/2$, while fixing $n=10^{5}$ and $p=4$.
-For the third experiment, we first fix $n=10^{5}$, $d=2d_{1}=2d_{2}=1000$ and $p=4$.
-Then, we vary $c$ equally spaced between $0.2$ and $1$ for Case 1, and we vary $\rho$ equally spaced between 00 and $0.9$ for Case 2.
+In the first experiment, we vary the sample size $n$ in $\{2,4,6\}\times 10^{4}$, while fixing $d=200$ and $p=4$.
+In the second experiment, we vary $d\in\{100,200,300,400\}$, while fixing $n=2\times 10^{4}$ and $p=4$.
+For the third experiment, we first fix $n=2\times 10^{4}$, $d=200$ and $p=4$.
+Then, we vary $c$ equally spaced between $0.2$ and $0.8$ for Case 1, and we vary $\rho$ equally spaced between $0.2$ and $0.8$ for Case 2.
 Figure 2: In the top panel, the error metric is plotted against varying $n$, $d$ and $c$ for Case 1. In the bottom panel, the error metric is plotted against varying $n$, $d$ and $\rho$ for Case 2.
-The error metrics $\text{Err}(\widehat{\mathbf{V}},\mathbf{V}^{\star})$, averaged over $500$ repetitions, are presented in Figure 1.
+Boxplots of the error metrics $\text{Err}(\widehat{\mathbf{U}},\mathbf{U}^{\star})$, averaged over $100$ repetitions, are presented in Figure 2.
 The results for CL and CL-GD, CLAIME and CLAIME-GD are very close to each other, as indicated by the proximity of their lines in Figures 2 (a)–(c).
@@
 This can be theoretically justified by their lower bounds being irrelevant to both $n$ and $d$, as shown in Theorem 3.6.
-From the top panel of Figure 1(c), it can be observed that all methods exhibit larger errors when $c$ decreases, or in other words, the SNR decreases.
-Comparing to the other methods, CLAIME and CLAIME-GD demonstrate superior performance in the small SNR regions where $c\leq 0.5$.
-Similarly, the bottom panel of Figure 1(c) shows a consistent increase in errors across all methods with higher values of $\rho$.
-CLAIME and CLAIME-GD again significantly outperform the other methods when $\rho$ is larger than $0.5$.
-This supports our initial motivation that the proposed method will be more robust in cases when the data is noisier.
+Moreover, it can be observed that KPCA behaves similarly to CL and Concate.
+From the top panel of Figure 2(c), we observe that all methods incur larger errors as $c$ decreases—that is, as the SNR worsens.
+Compared to the other methods, CLAIME and CLAIME-GD consistently demonstrate superior performance.
+This finding supports our initial motivation: the proposed method is more robust in noisier data settings.
+Similarly, in the bottom panel of Figure 2(c), while other methods exhibit only a slight decline in performance as $\rho$ decreases, CLAIME and CLAIME-GD again significantly outperform the alternatives across all values of $\rho$.
+We conducted additional simulation experiments in Supplementary S7.3.1 and S7.3.2 to empirically examine larger $d$ and account for model misspecification.
+Similar patterns were observed, demonstrating the robustness of our methods across broader settings.
+Furthermore, in Supplementary S7.3.3, we compared our approach against the Concate and CL methods with debiasing (Remark 3.2), and our method continued to outperform them.
 5 Application to Electronic Health Records Studies
-We consider training a joint representation of codified and narrative CUI features using summary EHR data of patients with at least 1 diagnostic code of rhematoid arthritis (RA) at Mass General Brigham (MGB).
-The MGB RA EHR cohort consist of 53,716 patients whose longitudinal EHR data have been summarized as co-occurrence counts of EHR concepts within a $30$-day window.
+We train a joint representation of codified and narrative CUI features using summary EHR data of patients with at least 1 diagnostic code of Rheumatoid Arthritis (RA) at Mass General Brigham (MGB).
+The MGB RA EHR cohort consists of 53,716 patients whose longitudinal EHR data have been summarized as co-occurrence counts of EHR concepts within a $30$-day window.
 We include all EHR diagnosis, medication, and procedure codes which have been rolled up to higher concept levels: diagnostic codes to PheCodes [^1] https://phewascatalog.org/phecodes, procedure codes to clinical classification system (CCS) [^2] https://www.hcup-us.ahrq.gov/toolssoftware/ccs_svcsproc/ccssvcproc.jsp, medication codes to ingredient level RxNorm codes (Liu et al., , 2005).
@@
 This results in $d=4,668$ features, including $d_{1}=3,477$ codified features with $1,776$ PheCodes for diseases, $238$ CCS codes for procedures, and $1,463$ RxNorm codes for medications, and $d_{2}=1,048$ CUI features.
-Based on the summary-level co-occurrence matrix, we can derive estimators for CLAIME, CL, and Concate, respectively.
+Based on the summary-level co-occurrence matrix, we can derive estimators for CLAIME, CL, Concate and KPCA, respectively.
 To assess the quality of the obtained embeddings, we utilize a benchmark previously introduced in Gan et al., (2025).
-Specifically, we conduct an evaluation focusing on the embeddings’ capability to identify established relationships among EHR concept pairs.
-Various categories of known relations, including similarity and relatedness, have been curated from online knowledge sources.
-Similar pairs of codified concepts were mainly generated based on code hierarchies such as the PheCode hierarchy.
-Similar pairs of CUIs were extracted from relationships within the online clinical database UMLS.
-Additionally, we assessed the similarity between mapped code-CUI pairs, leveraging UMLS to map codified concepts to CUIs.
-For relatedness among CUI-CUI pairs, we extracted information from the UMLS, considering major classes like “may treat or may prevent”, “classifies”, “differential diagnosis”, “method of”, and “causative”.
-Furthermore, related code-CUI pairs and code-code pairs were curated by mapping disease CUIs to PheCodes, drugs to RxNorm, and procedures to CCS categories.
-These mapped code pairs were then utilized to evaluate the system’s ability to detect relatedness among codified features or the relatedness between codified and NLP features.
-To assess different relationship types, we compute cosine similarities between embedding vectors of related pairs and randomly selected pairs.
-This approach enables the calculation of the Area Under the Curve (AUC) of the cosine similarities, providing a metric for distinguishing known pairs from random pairs.
-The random pairs selected have the same semantic type as the known pairs.
-For example, in investigating the relationship “may treat or may prevent,” we only consider disease-drug pairs.
-This evaluation strategy allows us to quantify and compare the performance of the EHR embeddings across various relationship types.
-Pairs Type Group CLAIME CL Concate Num
-CUI-Code Similar CUI-PheCode 0.920 0.920 0.920 345
-CUI-RXNORM 0.972 0.919 0.918 49
-summary 0.915 0.910 0.909 394
-Related May Treat (Prevent) 0.775 0.753 0.753 1302
-Classifies 0.905 0.895 0.895 744
-ddx 0.773 0.765 0.766 1169
-Causative 0.775 0.741 0.742 524
-summary 0.800 0.783 0.784 3739
-CUI-CUI Similar Parent 0.848 0.855 0.856 422
-Sibling 0.796 0.763 0.766 865
-summary 0.813 0.794 0.796 1287
-Related May Treat (Prevent) 0.767 0.765 0.766 257
-Classifies 0.936 0.924 0.925 45
-ddx 0.741 0.771 0.773 95
-Method_of 0.809 0.796 0.800 15
-summary 0.756 0.752 0.752 421
-Code-Code Similar PheCode Hierachy 0.914 0.909 0.909 3868
-Related ddx 0.790 0.796 0.795 5819
-Classifies 0.896 0.900 0.900 4525
-May Treat (Prevent) 0.734 0.727 0.725 4598
-Causative 0.754 0.756 0.754 2468
-summary 0.798 0.799 0.798 17410
-Table 1: AUCs of between-vector cosine similarity in detecting known similar or related pairs of codes, CUIs, and codes vs CUIs for embeddings trained by CLAIME, CL, and Concate methods.
-The last column indicates the number of known pairs.
-The results are presented in Table 1.
-It can be observed that CLAIME excels in embedding relationships across two distinct modalities: code and CUIs.
-In particular, for the similar CUI-RxNorm group, CLAIME achieves about $6\%$ gain in AUC compared to CL and Concate.
-It is also much better at capturing causative relationships between codes and CUIs, achieving a $4.4\%$ gain in AUC.
-Moreover, CLAIME is slightly better in representing the parent-child hierarchical relationships between NLP concepts, with a 2% gain in AUC compared to the other two methods.
+For multimodal learning, we are interested in our method’s ability to capture cross modality relationships.
+Specifically, we first assess the similarity between mapped code-CUI pairs, leveraging UMLS to map codified concepts to CUIs.
+For relatedness, we first curate UMLS relationships for CUI-CUI pairs, considering major classes like “may treat or may prevent”, “classifies”, “differential diagnosis”, “method of”, and “causative”.
+Then, we map disease CUIs to PheCodes, drugs to RxNorm, and procedures to CCS categories to obtain related code-CUI pairs.
+These mapped code pairs are then utilized to evaluate the system’s ability to detect relatedness between codified and NLP features.
+To assess different relationship types, we compute cosine similarities between embedding vectors of known and randomly selected pairs.
+This enables calculation of the Area Under the Curve (AUC), quantifying how well the embeddings distinguish known pairs from random ones.
+Random pairs are matched on semantic type; for instance, when evaluating the “may treat or may prevent” relationship, we consider only disease–drug pairs.
+This strategy allows us to compare embedding performance across relationship types.
+Type Group CLAIME CL Concate KPCA Number of known pairs
+Similar CUI-PheCode 0.910 0.911 0.911 0.916 345
+CUI-RxNorm 0.944 0.898 0.900 0.986 49
+Related May Treat (Prevent) 0.775 0.753 0.753 0.711 1302
+Classifies 0.905 0.895 0.895 0.887 744
+ddx 0.773 0.766 0.765 0.670 1169
+Causative 0.775 0.742 0.741 0.695 524
+Table 1: AUCs of between-vector cosine similarity in detecting known similar or related pairs of codes vs CUIs for embeddings trained by different methods.
+The results in Table 1 show that all methods are more effective at identifying similar pairs than related ones, as indicated by consistently higher AUCs in the “Similar” category.
+Among the methods, KPCA performs best for detecting similarity, achieving the highest AUCs for both CUI-PheCode and CUI-RxNorm pairs.
+However, when it comes to the more challenging task of detecting relatedness, the proposed CLAIME method consistently outperforms the baselines, especially KPCA.
+Notably, CLAIME shows a substantial advantage in identifying causative relationships, with an approximate 3.3% AUC improvement over CL and Concat, and an 8% gain over KPCA.
+These results underscore CLAIME’s robustness in capturing complex and nuanced associations between medical concepts.
 Since the data comes from the RA cohort, it is of interest for us to further perform case studies for RA-related codified and CUI concepts.
-We select $7$ key concepts for detailed analysis: rheumatoid arthritis, C-reactive proteins, stiffness of joint, Reiter’s disease, swollen joint, pneumococcal vaccine, and leflunomide.
+We select $7$ key concepts for detailed analysis: rheumatoid arthritis (RA), C-reactive proteins (CP), stiffness of joint (SoJ), Reiter’s disease (RD), swollen joint (SJ), pneumococcal vaccine (PV), and leflunomide (LF).
 For each of the $7$ concepts, we calculate its cosine similarity with all remaining concepts using the embeddings generated by CLAIME, Concate, and CL, respectively.
@@
 Concordance between similarity scores (from CLAIME, Concate, CL) and relevance scores (from GPT3.5 (OpenAI, , 2023), GPT4 (Achiam et al., , 2023) ) is assessed using Kendall’s tau rank correlation.
-Relevance scores are obtained by prompting GPT3.5 and GPT4 to rate concept relevance on a scale of 00 to $1$ between the given concept and its respective evaluation set.
-Figure 2: Feature-specific Kendall’s ranking correlation between pairs of CLAIME, Concate and CL with GPT3.5 and GPT4 feature scores for 7 concepts, namely, rheumatoid arthritis (RA), C-reactive proteins (CP), stiffness of joint (SoJ), Reiter’s disease (RD), swollen joint (SJ), pneumococcal vaccine (PV), and leflunomide (LF).
-Note that “GPTx” refers to GPT3.5 on the left and GPT4 on the right, and the lines for “CL-GPTx” overlap with “Concate-GPTx” and hence cannot be seen from the figure.
+Relevance scores are obtained by prompting GPT3.5 and GPT4 to rate concept relevance on a scale of 0 to $1$ between the given concept and its respective evaluation set.
+The detailed prompt is given in Supplementary S7.1.
+Figure 3: Feature-specific Kendall’s rank correlations between CLAIME, Concate, and CL with GPT-3.5 and GPT-4 scores for 7 concepts.
+‘GPTx’ denotes GPT-3.5 (left) and GPT-4 (right); note that CL-GPTx overlaps with Concate-GPTx and is not visible in the figure.
 The results, depicted in Figure 3, reveal that Concate and CL exhibit remarkably similar behavior, with the “CL-GPTx” lines obscured beneath “Concate-GPTx”.
@@
 Notably, negative rank correlations can be observed for CL-GPT3.5/4 and Concate-GPT3.5/4 regarding “leflunomide” and “pneumococcal vaccine” related codes and CUIs.
-To gain more insights, some detailed examples are provided in Table 2.
-ID Desc Concate CL CLAIME GPT3.5 GPT4
-Relatedness or similarity with Leflunomide
-PheCode:714.1 Rheumatoid arthritis 0.35 0.35 0.80 0.9 0.9
-PheCode:714.2 Juvenile rheumatoid arthritis 0.36 0.37 0.78 0.9 0.8
-C0242708 DMARDs 0.36 0.36 0.81 0.9 1.0
-RXNORM:5487 Hydrochlorothiazide 0.80 0.81 0.01 0.1 0.2
-RXNORM:6809 Metformin 0.79 0.79 0.01 0.2 0.2
-RXNORM:2556 Citalopram 0.78 0.78 0.05 0.1 0.1
-Relatedness or similarity with Pneumococcal vaccine
-C0151735 Injection site reaction NOS 0.29 0.29 0.74 0.8 0.8
-CCS:228 Prophylactic vaccinations and inoculations 0.29 0.29 0.64 0.9 1.0
-C1445860 Protein antibody 0.04 0.04 0.61 0.6 0.7
-RXNORM:10689 Tramadol 0.80 0.80 0.12 0.1 0.1
-RXNORM:2556 Citalopram 0.75 0.75 0.15 0.1 0.1
-RXNORM:11248 Cyanocobalamin 0.81 0.81 0.18 0.2 0.1
-Table 2: Cosine similarity between leflunomide, pneumococcal vaccine and some selected features.
-More specifically, leflunomide is a type of disease-modifying antirheumatic drugs (DMARDs) used to treat rheumatoid arthritis by reducing inflammation and permanent damage.
-Table 2 shows that both Concate and CL fail to capture the close connection between leflunomide, RA, and DMARDs.
-Additionally, they overestimate similarity with unrelated drugs like hydrochlorothiazide, metformin, and citalopram.
-In the case of pneumococcal vaccine, it is a vaccine targeting streptococcus pneumoniae infections and can potentially cause injection site reactions.
-It is related to protein antibodies as they all target the immune system (von Elten et al.,, 2014).
-Both Concate and CL struggle to capture this similarity and instead overestimate similarity with unrelated drugs like tramadol, cyanocobalamin, and citalopram, used for pain relief, vitamin B12 deficiency, and depressive disorders respectively.
+To gain more insights, some detailed examples are provided in Table S1 in Supplementary S7.1.
+As shown in Supplementary S6, CLAIME achieved strong performance when partitioning data elements into five modalities: diagnoses, procedures, medications, lab results, and CUIs.
+Its advantage is most evident in evaluating related relationships, with particularly pronounced improvements for ‘causative,’ ‘classifies,’ and ‘method of’ associations.
 6 Discussion and Conclusion
-In this paper, we propose the noisy multimodal log-linear production model for analyzing the multimodal EHR data and present the privacy-preserving algorithm CLAIME to estimate the multimodal feature embeddings with theoretical justification.
-Our algorithm is readily available for federated learning when multiple healthcare systems want to co-train their model but can not share patient-level data.
-It is also of great interest to consider a setting where different systems have overlapping but non-identical feature sets, in which case a block-wise matrix completion algorithm may be useful (Zhou et al.,, 2021).
-Although our current analysis focuses on two modalities, it can be extended to more than two modalities.
-The rich EHR data contains many different modalities such as genetic data and image data.
-It will be interesting to extend our current framework to accommodate these data.
-Besides, it will also be interesting to incorporate the patient demographic in the generative model to consider the effects of variables like sex and age.
-This paper focuses on a linear loss function, while non-linear loss functions such as mentioned in Remark 2.1 also deserve to be explored.
-Another interesting problem is the estimation of the patient embedding $\bm{c}_{i}$.
-The patient embedding is useful for many downstream tasks such as identifying “patients like me”.
+In this paper, we propose the noisy multimodal log-linear production model for analyzing multimodal EHR data and present the privacy-preserving algorithm CLAIME for estimating multimodal feature embeddings with theoretical justification.
+Our algorithm is readily applicable to federated learning when multiple healthcare systems wish to co-train their models without sharing patient-level data.
+It is also of great interest to consider settings where different systems have overlapping but non-identical feature sets, in which case block-wise matrix completion techniques may be useful (Zhou et al., , 2023).
+While our main development focuses on the two-modality setting, CLAIME is conceptually extensible to more than two modalities, as illustrated in Supplementary S6.
+Theoretical justification, however, remains challenging.
+A central insight of our work is that, in the two-modality case, the loss function admits a closed-form global minimizer directly connected to the SVD of the PMI matrix.
+This reduction is crucial: it allows us to analyze the performance of the CLAIME representation through the top-$p$ singular space of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$.
+By contrast, the straightforward multimodal extension lacks such an explicit form, making the global minimizer analytically intractable.
+This limitation reflects a broader challenge in the theory of contrastive learning, where rigorous analyses remain largely confined to the two-view case (Nakada et al., , 2023; Dufumier et al., , 2024).
+Moreover, our current extension assumes access only to pairwise co-occurrence matrices (Finlayson et al., , 2014; Beam et al., , 2020).
+If patient-level data were available, one could, in principle, construct higher-order co-occurrence tensors that capture joint dependencies across three or more modalities.
+Tensor SVD or related decompositions could then provide a natural way to generalize CLAIME embeddings with accompanying theoretical guarantees.
+Such a tensor-based framework could reveal interactions that are invisible to pairwise models, for example uncovering clinical pathways that involve coordinated use of lab tests, diagnoses, and treatments.
+Developing the statistical theory and scalable algorithms for such tensor-based generalizations remains an important direction for future work.
+Beyond the linear loss considered here, exploring nonlinear losses such as those in Remark 2.4 is a natural next step.
+While a linear inner-product model captures only additive associations between modalities, nonlinear mappings (e.g., replacing the bilinear form $\mathbf{V}_{1}\mathbf{V}_{2}^{\scriptscriptstyle\sf T}$ with a neural network) could model richer dependencies.
+However, ensuring identifiability and statistical guarantees would require new theory beyond classical low-rank approximation, making the balance between flexibility, interpretability, and tractability a key open challenge.
+Another promising direction is the estimation of patient embeddings $\bm{c}_{i}$, which could enable clustering into clinically meaningful subgroups, modeling disease trajectories, and powering personalized applications such as “patients like me” recommendations.
+Such embeddings may be derived through joint factorization of patient–feature co-occurrence matrices or by incorporating temporal information from longitudinal data, with a theoretical framework for their statistical properties—consistency, stability, and generalization—being essential to ensure reliability in practice.